### v2

# GraphTeam: Facilitating Large Language Model-based Graph Analysis via Multi-Agent Collaboration

Abstract.
Graphs are widely used for modeling relational data in real-world scenarios, such as social networks and urban computing. While large language models (LLMs) have achieved strong performance in many areas, existing LLM-based graph analysis approaches either integrate graph neural networks (GNNs) for specific machine learning tasks (e.g., node classification), limiting their transferability, or rely solely on LLMs’ internal reasoning ability, resulting in suboptimal performance. To address these limitations, we take advantage of recent advances in LLM-based agents, which have shown capabilities of utilizing external knowledge or tools for problem solving. By simulating human problem-solving strategies such as analogy and collaboration, we propose a multi-agent system based on LLMs named GraphTeam, for graph analysis. GraphTeam consists of five LLM-based agents from three modules, and the agents with different specialities can collaborate with each other to address complex problems. Specifically, (1) input-output normalization module: the question agent extracts and refines four key arguments (e.g., graph type and output format) from the original question, facilitating the problem understanding, and the answer agent organizes the results to meet the output requirement; (2) external knowledge retrieval module: we first build a knowledge base consisting of relevant documentation and experience information, and then the search agent retrieves the most relevant entries from the knowledge base for each question. (3) problem-solving module: given the retrieved information from search agent, the coding agent uses established algorithms via programming to generate solutions, and in case the coding agent does not work, the reasoning agent will directly compute the results without programming. Extensive experiments on six graph analysis benchmarks demonstrate that GraphTeam achieves state-of-the-art performance with an average 25.85% improvement over the best baseline in terms of accuracy. The code and data are available at https://github.com/BUPT-GAMMA/GraphTeam.
Large Language Model, Multi-agent System, Graph Analysis
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Graph is a commonly used data structure for modeling relational data. Correspondingly, tasks and methods of graph analysis have been extensively exploited in AI field (Brown et al., 2020; Team et al., 2024; Anthropic, 2024) and achieved significantly success in many fields, such as social network (Myers et al., 2014; Min et al., 2021) and unban computing (Yan et al., 2024; Zou et al., 2024). Recently, with the surge of LLM, some researchers begin to explore the feasibility of graph analysis with LLMs (Tang et al., 2024; Zhang et al., 2024a; Ouyang et al., 2024; Sanford et al., 2024).
Existing LLM-based graph analysis methods mainly fall into two categories. The first category (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023) usually combines LLMs with graph neural networks (GNNs) for modeling. A typical paradigm is to first encode the input graph with a GNN, and then feed the node encodings to an LLM as language token embeddings. However, this line of work is specialized for graph machine learning tasks, especially node classification, and has limited transferability to general graph analysis problems. The second category (Fatemi et al., 2023; Chen et al., 2024b) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c). Strategies like in-context learning (Brown et al., 2020) and chain-of-thought reasoning (Wei et al., 2023) are widely used in these methods to improve performance. However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents (Xi et al., 2023; Wang et al., 2024b). Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems (Chen et al., 2023; Qian et al., 2024; Hong et al., 2023; Wu et al., 2023), where agents with different specialities can discuss and collaborate with each other to address complex problems. However, previous multi-agent frameworks are designed for general purposes and have few knowledge relevant to graph analysis, and thus fail to outperform the aforementioned methods tailored for graph analysis.
Figure 1. The overall pipeline of our multi-agent system GraphTeam (left), and the comparison between GraphTeam and state-of-the-art baseline on six benchmarks (right).
To overcome the limitations of previous work, we aim to design a multi-agent system based on LLMs with its design and implementation specialized for graph analysis. Inspired by the problem-solving process of humans (Robson and Pólya, 1945; Ramnarine-Rieks, 2011; Bybee et al., 2006; Brand-Gruwel et al., 2009; Savery and Duffy, 1995; Eisenberg and Berkowitz, 1990), we outline our system GraphTeam with three core modules: an input-output normalization module, an external knowledge retrieval module, and a problem-solving module. The overall pipeline of GraphTeam is shown in Fig. 1. (1) During human problem-solving processes, clearly understanding and formalizing the problem is the first step (Bybee et al., 2006; Robson and Pólya, 1945), and correctly answering the question in required formats is also important. Therefore, we introduce the input-output normalization module with two different LLM-based agents: the question agent aims to extract and refine four key arguments (e.g., graph type and output format) from the original question description, facilitating the system to better understand the problem; the answer agent aims to reformat the results to meet the output requirement. (2) During the solving process of a hard problem, humans will seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to previously solved ones (Bybee et al., 2006). Therefore, we first build a knowledge base consisting of relevant documentation and experience information. Then we introduce the external knowledge retrieval module with a search agent to retrieve the most relevant entries from the knowledge base for each question. The retrieved information is expected to benefit the downstream solving process. (3) In the problem-solving module, we introduce a coding agent that uses established algorithms via programming to generate solutions. If the coding agent fails to generate effective codes, a reasoning agent will directly infer the results without programming. We conduct extensive experiments on six graph analysis benchmarks, including Talk like a Graph (Fatemi et al., 2023), LLM4DyG (Zhang et al., 2024b), GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), Graph Instruct (Luo et al., 2024) and a newly introduced GNN-AutoGL based on AutoGL (Guan et al., 2021). These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs. The experimental results and ablation studies demonstrate the effectiveness of GraphTeam as well as the necessity of each module. In summary, our contributions are three-fold:
$\bullet$ Different from previous paradigms that utilize LLMs for graph analysis, we propose to facilitate the analysis via the collaboration of multiple LLM-based agents.
$\bullet$ Inspired by the problem-solving process of humans, we propose a carefully designed framework GraphTeam with five agents from three functional groups, i.e., input-output normalization, external knowledge retrieval, and problem-solving modules. The agents can work in cooperation with a due division of labor.
$\bullet$ We conduct extensive experiments over very recent graph analysis benchmarks. Experiments show that GraphTeam achieves state-of-the-art (SOTA) performance on all six benchmarks, with an average 25.85% improvement on accuracy over the best baseline.
2. Related Work
2.1. LLM-based Graph Analysis
2.1.1. Benchmarks
To evaluate the ability of large language models (LLMs) in solving graph analysis problems, researchers have developed various benchmarks (Das et al., 2024; Chen et al., 2024a; Guo et al., 2023; Fatemi et al., 2023). Among them, NLGraph (Wang et al., 2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations. Graphwiz (Chen et al., 2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete problems. GraphInstruct (Luo et al., 2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems. Talk like a Graph (Fatemi et al., 2023) measures basic graph understanding abilities, while LLM4DyG (Zhang et al., 2024b) focuses on dynamic graph tasks related to temporal and spatial information. In our experiments, we employ the above five benchmarks and also build an additional one focusing on GNNs for assessing the abilities of different methods.
2.1.2. Methods
There is a rising trend of research to utilize LLMs for various graph analysis tasks (Li et al., 2024b; Jin et al., 2024). Here we focus on the most relevant methods that employ LLMs as predictors (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023). Related work can be roughly divided into two categories. The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification (Tang et al., 2024; Zhang et al., 2024a). For example, GraphGPT (Tang et al., 2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification. GraphTranslator (Zhang et al., 2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification. However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse analysis tasks in the above benchmarks. The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning (Brown et al., 2020), chain-of-thought reasoning (Wei et al., 2023) for inference (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c). For example, GUNDAM (Ouyang et al., 2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference. However, recent studies have shown that the reasoning depths of current LLMs are still shallow (Lam et al., 2024; Li et al., 2024a), and will probably fail for graph analysis problems that require complex reasoning. In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
2.2. Multi-Agent Collaboration of LLMs
Autonomous agents aim to solve problems with self-directed planning and actions. Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions (Xi et al., 2023; Bang et al., 2023; Li et al., 2023; Wang et al., 2024b). Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where multiple agents work in cooperation with a due division of labor.
For example, ChatDev (Qian et al., 2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration. MetaGPT (Hong et al., 2023) is a meta-programming framework that simulates human workflows. With proper developing pipeline, MetaGPT can generate effective programming solutions. AgentVerse (Chen et al., 2023) is a multi-agent framework inspired by human collaboration dynamics. It involves four iterative stages of expert recruitment, decision discussion, action and evaluation. AutoGen (Wu et al., 2023) is another multi-agent framework for general purposes. Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks. MACRec (Wang et al., 2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration. This system includes various specialized agents and can easily solve various recommendation tasks. However, these multi-agent frameworks are not specialized for graph analysis, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems. Consequently, they can hardly beat the SOTA performance on the aforementioned graph analysis benchmarks.
3. Methodology
In this section, we propose GraphTeam, a multi-agent system based on LLMs for graph analysis. We will start by introducing the overall framework, and then present the details of each agent in the system.
3.1. Framework Overview
As shown in Fig. 2, the entire framework of GraphTeam consists of five agents from three functional groups, i.e., question and answer agents for input-output normalization, search agent for external knowledge retrieval, as well as coding and reasoning agents for problem solving.
Figure 2. The overall framework of GraphTeam, which includes five agents from three functional groups.
Input-Output Normalization: This functional group aims to refine key information from input questions, and map the results to desired output formats. Specifically, the question agent is responsible for processing the original problem descriptions, extracting key arguments to assist the system in solving it. The answer agent is responsible for transferring the formats of calculated results, and conduct self-checking to ensure the consistency with problem requirements.
External Knowledge Retrieval: This functional group only includes the search agent to extract relevant information from external knowledge base. Specifically, we first build a knowledge base for graph analysis based on documentation of popular Python libraries and experiences of previous problem-solving processes. With problems refined by the upstream question agent as queries, the search agent then retrieves information beneficial to downstream problem solving from the knowledge base.
Problem Solving: This functional group aims to calculate the results based on the refined problem of question agent and the extracted knowledge of search agent. Specifically, the coding agent attempts to write Python codes to solve the problem. When the generated codes cannot be executed normally, a retry mechanism will be employed to fix the errors in the codes. If the coding agent still fails after several retries, the reasoning agent will solve the problem directly without programming.
3.2. Question Agent
Abstraction is a preliminary step for human beings to solve problems (Robson and Pólya, 1945). Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions. Formally, given a graph analysis problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
(1) $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q),$
where $q_{r}$ is a condensed version of original instruction, $q_{t}$ indicates the graph type (e.g., directed, undirected, weighted, etc.), $q_{g}$ contains the graph to be analyzed, and $q_{f}$ specifies how the answer should be presented (e.g., numerical value, list, etc.). By explicitly identifying these key information, downstream agents can better understand a task and its requirements, thereby enhancing the overall analysis capability. An example of the question agent is shown below:
Original Question
In an undirected graph … The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5).
Q: Is there a path in this graph that visits every node exactly once? …
A: In the answer, …, Output format such as: … 1, 4, 8.
Reformatted Question
refined_question: Is there a path in this graph that visits …
graph_type: undirected
input_data: ’Nodes’: [0, 1, 2, 3, 4, 5, 6],
’Edges’: [[0, 3], [0, 1], [1, 4], [1, 6], [2, 4], [3, 5], [3, 6], [4, 5]]
output_format: Yes. The path can be: 1,4,8.
3.3. Search Agent
During the solving process of a hard problem, it is common for human beings to seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to a previously solved one (Bybee et al., 2006). Therefore, we first build a knowledge base with documentation and experience information to enhance the programming ability of the multi-agent system, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
3.3.1. Knowledge Base Construction
The knowledge base has two parts, i.e., documentation base $\mathcal{K}_{doc}$ and experience base $\mathcal{K}_{exp}$. Examples of documentation and experience entries are shown below:
Documentation Information
api_name: hamiltonian_path
description: Returns a Hamiltonian path in the given …
parameters: ”G : NetworkX graph” , …
returns: ”path : list”: ”A list of nodes which form a Hamiltonian path in G.”
examples:
    import NetworkX as nx
    nx.tournament.hamiltonian_path(G)
Documentation. We crawl the documentation of relevant Python libraries from the Internet, and each entry in the documentation corresponds to a specific application interface (API). Specifically, each entry includes fields such as API name, description, parameters, returns, and examples.
Experience. We use the system without experience base as the solver, and apply it on the training set to collect experiences. Note that each problem in the training set $\mathcal{D}_{\text{train}}$ is usually associated with problem type $t\in\mathcal{T}$. Thus we collect candidate experiences $\mathcal{K}_{exp}^{t}$ for each problem type $t$, and evaluate the utility of each experience on the validation set $\mathcal{D}_{\text{valid}}$. Only the experience with the highest utility will be kept in each $\mathcal{K}_{exp}^{t}$.
[…]
Experience Information
question: In an undirected graph, …, The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5) Q: Is there a path in this graph …
answer: Yes. The path can be: 0, 1, 6, 3, 5, 4, 2.
meta information:
thought: The experience from this task involved solving the problem of finding a Hamiltonian path. …
code:
    import NetworkX as nx
    def hamiltonian_path_backtrack(graph, path):
3.3.2. Knowledge Retrieval
For effective information retrieval, we build vector databases with LlamaIndex (lla, 2024) for documentation and experiences, respectively. Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database. Note that the experiences of similar problems contains detailed solutions, and are usually more helpful for downstream agents than documentations. Therefore, we first use similarity matching via LlamaIndex to score the experiences. If the largest similarity score exceeds a predefined threshold $\delta$, the corresponding experience will be directly returned by the search agent without querying related documentation information. Otherwise, the search agent will return the most related entries in documentation base. Finally, either the experience or the documentation information returned by the search agent will be passed to the downstream agents for retrieval-augmented generation (RAG). For simplicity, we denote the operation of the search agent as
(2) $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta).$
Algorithm 1 Problem-Solving Experience Collection
1:Training dataset $\mathcal{D}_{\text{train}}$, validation dataset $\mathcal{D}_{\text{valid}}$, candidate experience size $N_{exp}$ for each problem type $t\in\mathcal{T}$.
2:Experience knowledge base $\mathcal{K}_{exp}$.
3:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{train}}$ do
4:     if $|\mathcal{K}_{exp}^{t}|$ $N_{exp}$ then
[…]
6:         if answer equals to ground truth $a$ then
[…]
8:         end if
9:     end if
10:end for
11:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{valid}}$ do
12:     for $e\in\mathcal{K}_{exp}^{t}$ do
[…]
14:         if answer equals to ground truth $a$ then
15:              $\text{Utility}_{e}\leftarrow\text{Utility}_{e}+1$
16:         end if
17:     end for
18:end for
19:for $t\in\mathcal{T}$ do
20:     $\mathcal{K}_{exp}^{t}\leftarrow\text{argmax}_{e\in\mathcal{K}_{exp}^{t}}\text{Utility}_{e}$
21:end for
22:$\mathcal{K}_{exp}\leftarrow\{\mathcal{K}_{exp}^{t}\}_{t\in\mathcal{T}}$
3.4. Coding and Reasoning Agents
Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems. In practice, the generated codes may encounter compilation or running errors. Following the trial-and-error strategy (Robson and Pólya, 1945; Ramnarine-Rieks, 2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages. More formally, the operation of coding agent in the $n$-th trial can be written as:
(3) $\begin{split}(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},\text{knowledge}_{q},\\ \{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1}),\end{split}$
where $\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n}$ are respectively the execution result, generated code, and error message of the $n$-th trial. The last two parameters are the codes and error messages from previous $n-1$ trials. If the codes runs normally, we define $\text{error}_{q}^{n}$ as $\tt None$.
If the coding agent fails to fix the codes before reaching the maximum number of trials, we will employ the reasoning agent to directly answer the question without programming:
(4) $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f}).$
3.5. Answer Agent
Based on the output format $q_{f}$, the answer agent will organize the results given by the coding or reasoning agents as final results. Similar to the retry mechanism of coding agent, we also introduce a self-checking mechanism that iteratively refines the results according to format requirement $q_{f}$. Formally, the operation of answer agent can be written as:
(5) $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1}),$
where $\text{output}_{q}^{n}$ is the output in the $n$-th iteration, and $\text{output}_{q}^{0}$ is $\text{result}_{q}$ generated by the coding or reasoning agents.
Input
result: 0, 1, 6, 3, 5, 4, 2.
output_format: Yes. The path can be: 1, 4, 8.
Answer
result: Yes. The path can be: 0, 1, 6, 3, 5, 4, 2.
3.6. Summary
We formalize the inference pipeline of GraphTeam in Alg. 2. Most agent functions (including Question, Coding, Reasoning and Answer) in Alg. 2 are all implemented by an LLM with different prompt templates. The Search function is implemented based on LlamaIndex (lla, 2024), and the Solver function in Alg. 1 is implemented by a simplified version of GraphTeam without experience base. We will provide more implementation details in the appendix.
Algorithm 2 Inference Pipeline of GraphTeam
1:Problem described by question $q$, knowledge base $\mathcal{K}$, threshold $\delta$ for determining whether experience or documentation is retrieved, maximum number of retries $N_{retry}$, number of self-checking iterations $N_{check}$.
2:Final answer of the problem $\text{output}_{q}$.
3:$(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
4:$\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
5:$n=0$
6:while $\text{error}_{q}^{n}\neq\tt{None}$ and $n<N_{retry}$ do
7:     ++n
8:     $(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},$ $\text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1})$
9:     $\text{result}_{q}\leftarrow\text{result}_{q}^{n}$
10:end while
11:if $\text{error}_{q}^{n}\neq\tt{None}$ then
12:     $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f})$
13:end if
14:$\text{output}_{q}^{0}\leftarrow\text{result}_{q}$
15:for $n=1,2,\dots,N_{check}$ do
16:     $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1})$
17:end for
4. Experiments
In this section, we conduct experimental evaluation to answer the following research questions (RQs): RQ1: How effective is our proposed GraphTeam compared to the state-of-the-art (SOTA) baselines? RQ2: Have each component of GraphTeam played their roles effectively? RQ3: How does our framework perform with respect to different task categories and output formats of graph analysis problems? RQ4: How do the hyper-parameters of certain components in the system have an impact on the overall performance?
4.1. Experimental Setup
4.1.1. Datasets
To evaluate the effectiveness of our system in solving graph analysis problems, we use six benchmarks for LLM-based graph analysis: Talk like a Graph (Fatemi et al., 2023) primarily studies some fundamental problems in graph analysis and explores the cognitive abilities of LLMs regarding graphs. LLM4DyG (Zhang et al., 2024b) focuses on basic problems related to dynamic graphs, while GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), and Graph Instruct (Luo et al., 2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory. GNN-AutoGL (details shown in Appendix B) is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL (Guan et al., 2021).
4.1.2. Metrics
For benchmarks with definite answers (Fatemi et al., 2023; Chen et al., 2024a; Wang et al., 2024a; Luo et al., 2024; Zhang et al., 2024b), we use accuracy as the evaluation metric. For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria.
Specifically, we extract the generated code from the model response and compare the execution result with the reference solution to assess whether the model-generated results match the given standard answers or reference solutions. For the first five benchmarks (Fatemi et al., 2023; Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024; Wang et al., 2024a), the evaluation results are binary based on the exact match between the execution results and the standard answers or reference solutions. For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0. Otherwise, we check whether the hyper-parameters(eg. model name, dataset name, hidden dimension, dropout) meet the requirements. Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
4.1.3. Settings
We choose GPT-4o-mini (OpenAI, 2024) as the base LLM for all agents of our GraphTeam. For Python libraries, we use NetworkX version 3.3111https://networkx.org/ and AutoGL version 0.4.0222https://mn.cs.tsinghua.edu.cn/autogl/documentation/, together with Python version 3.10.14. The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$. The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
4.1.4. Baselines
As the SOTA methods for each benchmark are quite diverse and some of which are specialized for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM (Ouyang et al., 2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers (Sanford et al., 2024) as the SOTA; for GraphWiz, LLM4DyG, and GraphInstruct (Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024), we select the best model in their benchmark papers. For GraphWiz (Chen et al., 2024a), we did not choose GraphAgent-Reasoner (Hu et al., 2024) because it only provide results for 6 out of 9 categories.
4.2. Main Results (RQ1)
Table 1. Performance comparison on six graph analysis benchmarks in terms of accuracy (%).
Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
Main GPT-4o-mini 71.67 47.61 51.40 58.03 53.67 3.28
SOTA 97.30 65.00 79.70 53.71 43.53 -
GraphTeam 99.13 88.61 97.80 94.81 88.14 48.44
     - question agent 85.87 62.81 79.22 59.18 85.86 14.53
Input-Output Normalization      - answer agent 94.80 87.97 96.70 94.27 85.25 45.78
     - documentation 92.03 87.97 96.40 94.07 86.12 46.88
Knowledge Retrieval      - experience 89.96 82.83 54.70 60.26 76.64 4.76
     - coding agent 69.40 46.39 39.30 59.33 61.09 -
     - retry 91.90 83.53 87.70 89.33 85.66 44.22
Problem Solving      - reasoning agent 92.77 54.17 96.90 94.00 85.45 46.36
Table 2. Performance comparison on the hard subset of six graph analysis benchmarks in terms of accuracy (%).
Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
Main GPT-4o-mini 47.90 28.92 44.93 42.56 23.19 3.28
SOTA 93.90 70.50 44.96 33.90 11.71 -
GraphTeam 99.20 76.25 98.14 98.33 73.33 48.44
     - question agent 70.00 39.75 69.82 53.67 73.29 14.53
Input-Output Normalization      - answer agent 97.40 75.25 97.51 96.33 70.00 45.78
     - documentation 98.90 75.25 97.51 94.33 69.76 46.88
Knowledge Retrieval      - experience 98.40 62.92 42.32 70.56 63.95 4.76
     - coding agent 64.00 38.33 19.21 51.56 38.24 -
     - retry 95.40 61.00 40.27 85.44 70.95 44.22
Problem Solving      - reasoning agent 98.90 66.08 97.51 95.67 69.62 46.36
We present the results in Tab.1. GraphTeam surpasses previous SOTA by 1.83-44.61% in accuracy, and outperforms the base LLM (GPT-4o-mini) by 27.46-46.40% in accuracy. Such significant improvements fully demonstrate the effectiveness of our carefully designed multi-agent system. Besides, the time cost for answering each problem is less than 3 seconds, and the monetary cost of API usage is less than \$0.01.
To further demonstrate GraphTeam’s performance on challenging problems, we assess its performance on the “hard” subset of each benchmark. For GraphWiz, we adhere to the official partition and report results on their designated hard set. For other benchmarks, which all have their problems categorized into different types or domains, we employ a model-based approach: we identify the three categories where previous SOTA performs poorest and select problems from these categories. This selection method ensures we target the most challenging aspects of each benchmark.
Tab. 2 summarizes the system’s performance on these high-difficulty questions. Notably, GraphTeam maintains high accuracy on the Talk like a Graph hard set (99.20%), and shows significant improvements over SOTA for NLGraph (98.14% v.s. 44.96%) and LLM4DyG (98.33% v.s. 33.90%). While performance drops are observed for GraphWiz and GraphInstruct compared to the general set, GraphTeam still substantially outperforms both GPT-4o-mini and SOTA methods. This consistent and strong performance across diverse hard sets underscores GraphTeam’s robustness and its ability to effectively handle complex graph-related tasks.
4.3. Ablation Study (RQ2)
To evaluate the contribution of each component, we conduct a systematic ablation study. We assess the impact of Input-Output Normalization by removing the question or answer agent. We examine the influence of Knowledge Retrieval components in the search agent, specifically documentation and experience. Additionally, the effectiveness of Problem Solving designs are evaluated by eliminating the coding or reasoning agent, as well as the retry mechanism from the coding agent. This comprehensive ablation enables us to quantify each component’s significance to the system performance. The overall results are presented in Tab.1 and 2.
Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.44%) by breaking the problem into individual components. For a few benchmark such as GraphInstruct (Luo et al., 2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.28%). The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.54%- 3.33%).
External Knowledge Retrieval: The search agent plays a significant role. When solving graph analysis problems, providing the large language model with prior knowledge of related problem-solving experiences or API documentation can help the model solve the problem more effectively. The impact of experience is greater than that of documentation, indicating that LLMs solve problems in a way similar to humans with an average improvement of 24.63%. They can derive better solutions from prior experiences, effectively addressing similar problems. Documentation can enhance the model’s programming capabilities. When relevant experiences are not available, documentation is still needed to learn graph analysis methods and subsequently solve corresponding problems (0.64% to 7.1%).
Problem Solving: The coding agent plays a crucial role. Without programming and relying solely on reasoning, the system’s performance would significantly decrease, indicating that programming is an effective approach for solving graph analysis problems (27.05% to 58.50%). The retry mechanism does not always have a significant overall effect, but it shows clear benefits for hard problems with a maximum improvement of 57.87%. This indicates that Hard problems require more complex code, which cannot be written correctly in a single attempt and needs multiple tries to solve effectively. The reasoning agent does not have a significant impact on problems where programming achieves good results, but it plays an important role in problems where programming is insufficient and reasoning is required (0.81% to 34.44%).
The contribution of each component of the system varies across different benchmarks. Since the core problem-solving component of this system is the coding agent, programming to solve problems is a crucial and effective approach for graph analysis. For GNNs, in particular, coding is an indispensable component. Therefore, removing the coding agent leads to a significant drop in performance for all benchmarks (42.24% in average). Some benchmarks do not have strict requirements for the answer format (NLGraph, GraphWiz, and LLM4DyG), so removing the answer agent does not significantly impact the overall system performance (0.76% in average). Since Talk like a Graph contains simple questions, most problems can be solved correctly through coding without the need for the reasoning agent, hence the benchmark has minimal need for the reasoning agent (6.36%). Since GNN-AutoGL lacks documentation information, this module does not affect system performance (1.56%). Similarly, GraphInstruct’s problem-solving does not rely on documentation, so documentation has little impact on the system’s performance on this benchmark (0.04%).
Figure 3. Performance with respect to different task categories.
Figure 4. Performance with respect to different output formats.
Figure 5. Hyper-parameter analysis of four hyper-parameter in the proposed GraphTeam.
4.4. Analysis on Different Groupings (RQ3)
To provide a more nuanced understanding of GraphTeam’s capabilities and limitations, we conduct a detailed analysis of its performance across various task categories and output formats. This fine-grained approach allows us to identify specific strengths and weaknesses.
Group by Task Categories. We propose a categorization scheme that covers a broad spectrum of graph-related problems. We classify the questions into four categories based on the nature of the task: (1) basic: fundamental graph understanding, (2) algorithm-macro-level: analysis of entire graph structures, (3) algorithm-micro-level: local properties of nodes and edges, and (4) GNN: graph neural networks. As shown in Fig. 3, GraphTeam achieves an accuracy of over 90% on both basic and macro-level algorithms, and 85.2% on micro-level tasks, indicating that the system can effectively handle common graph analysis tasks. However, all systems, including GraphTeam, struggle with GNN-related tasks, suggesting a need for improvement in this area.
These results reveal interesting insights: For micro-level tasks, removing the reasoning agent causes a significant 21.2% drop in performance, highlighting the complexity of these problems and the limitations of the coding agent alone. In GNN-related tasks, while the coding agent and experience are important, the question agent plays a crucial role (38.8% impact when removed). This underscores the importance of effectively extracting key information from GNN problem statements. These findings point to areas for focused improvement, particularly in enhancing reasoning capabilities for micro-level tasks and information extraction for GNN-related problems.
Group by Output Formats. To account for the diversity in problem outputs, we categorize the questions based on their required response formats: (1) yes/no, (2) digits, (3) list/set, and (4) others. This classification allows us to assess system performance across different types of graph analysis outputs. As shown in Fig. 4, GraphTeam demonstrates strong performance on yes/no, digits, and list/set outputs, achieving an accuracy of over 85% for each of these categories. However, the system’s performance drops notably for more complex output formats, with an accuracy of only 70.1%. This decline in accuracy for the “others” format indicates that GraphTeam, particularly its answer agent, needs enhancement in handling advanced data structures and more complex output formats. Improving the system’s capability to generate and structure complex outputs could significantly boost its overall effectiveness in addressing a wider range of graph analysis problems.
4.5. Hyper-parameter Sensitivity (RQ4)
We conduct a study on key hyper-parameters using the NLGraph (Wang et al., 2024a) benchmark to determine optimal settings or optimization strategies. Fig. 5 presents our analysis of several important hyper-parameters. For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance. The similarity matching threshold $\delta$ between experiences and documents achieves optimal problem-solving performance at $0.85$, enabling correct matching of similar experiences when available and document retrieval when experiences are absent. Increasing the maximum retries of the coding agent and the number of self-checking iterations of the answer agent improves system performance, but with diminishing returns as these values increase, eventually plateauing. These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
5. Conclusion
To address the limitations of previous LLM-based graph analysis methods, we draw inspiration from human problem-solving processes, and design a multi-agent system named GraphTeam based on LLMs. GraphTeam consists of five agents from three key modules, namely input-output normalization, external knowledge retrieval, and problem solving. The agents can cooperate with each other for solving complex graph analysis problems. Extensive experiments on six benchmarks demonstrate that GraphTeam achieves state-of-the-art results with an average improvement of 25.85% in terms of accuracy. For future work, we will consider to fine-tune the LLMs in GraphTeam for better performance, and further optimize the system pipeline to accommodate more tasks.
Appendix A Prompt Templates
A.1. Question Agent
Requirement Analyst
Here is the task: task.
As the assistant_role, you need to extract key information from the task, specifically focusing on the following elements:
Graph_Type(str): This refers to the type of graph mentioned in the task, such as directed, undirected, weighted, dynamic, etc. If the task does not specify a graph type, it should be set to undirected.
Input(str): This refers to the graph data or graph file required by the task. You must extract **all relevant information** about the graph data, including but not limited to graph representation (adjacency list, adjacency matrix, edge list, etc.), node and edge details, and any other relevant information.
Problem(str): This is the specific problem that needs to be solved.
Output_Format(str): This is the format of the output answer mentioned in the task. If the task does not specify a format, it should be set to None.
Note that we must ONLY discuss these key elements and do not discuss anything else! Please provide the extracted information in JSON format. Your answer must be in JSON format.
A.2. Search Agent
Research Assistant
You are tasked as a assistant_role to retrieve relevant information for the task: query.
Use the available knowledge base and memory to find the most accurate and relevant information related to the task.
If the query is too ambiguous, try to clarify it using previous context.
Please provide the results in a structured format.
A.3. Coding Agent
Graph Learning Specialist
Based on the available search result: search_result, your goal is to write Python code that solves the specified graph learning problem using libraries like NetworkX.
If there is no search result, generate the code based on your expertise.
ATTENTION: This is the actual problem that needs to be solved, based on the provided information. Please disregard the example above while solving this real task.
Graph_Type: Graph_Type
Input: input,
Problem: problem,
Output Format: output_format.
The output code must adhere to the following format:
    ‘‘‘python
to ensure easy extraction.
If the code execution failed. Please analyze the error message and the code provided to identify the issuescausing the failure in code execution. Based on your analysis, revise the code and provide a corrected version that should execute successfully.
Here is the error message: error_message
The revised code must be formatted as follows for easy extraction:
    ‘‘‘python
A.4. Reasoning Agent
Graph Learning Expert
As a assistant_role, your task is to analyze and solve a graph analysis problem.
Despite any challenges that previous attempts may have encountered, your goal is to derive a solution leveraging your expertise in graph theory and machine learning.
Here is the task based on the extracted information: Input: input, Problem: problem, Output Format: output_format.
Provide a clear, concise answer and present your solution in the following format for easy extraction:
’###output:
solution
A.5. Answer Agent
Output Format Specialist
As a assistant_role, your role is to ensure that the generated output strictly adheres to the specified format requirements.
You will be provided with both the required output format (example format) and the actual output. Your task is to compare the two and determine if the actual output matches the format requirements. If the output does not meet the format, you must adjust the format accordingly while **preserving every part of the original content, meaning, and intent without deleting or changing any sections**.
Important: The required format provided is an **example** that illustrates the expected structure, not the actual content or final answer. Do **not** alter the original meaning, sections, sentences, or intent of the actual output. Your task is only to adjust the format if necessary.
When adjusting the format, ensure that **no parts are removed, shortened, or reworded in a way that changes the original meaning or omits content**. The adjustment should focus on structural changes such as spacing, indentation, or organization.
Once the adjustment is made, compare the adjusted output with the original to validate that **all sections, sentences, and words are fully preserved** in the final output.
Here are the provided details:
- **Required format (example)**: output_format
- **Actual output** (content to be reviewed): output
Carefully review both and ensure that the actual output conforms to the structure of the example format, while **maintaining the original content, meaning, sections, and intent**.
Note: Only adjust the **format** if necessary. Do **not** change the actual content, meaning, or intent of the output. If the format is already correct, no adjustment is needed.
Your response should be in JSON format with:
- **need_adjustment** (True/False): Indicating whether an adjustment is required.
- **output** (str): If no adjustment is needed, this should match the original output exactly. If an adjustment is needed, this should contain the adjusted output, with only format changes applied, without altering the content, meaning, sections, or intent.
Appendix B GNN-AutoGL Benchmark Construction
For the GNN-AutoGL benchmark, we use a unified template for construction. Specifically, we ask the large language model (LLM) to write code that constructs a GNN model using AutoGL based on the given prompt. Given the numerous model parameters, specifying each parameter and requiring the LLM to write corresponding model code is challenging. Therefore, we select only some parameters as requirements and evaluate whether the LLM correctly includes them. For a problem with $K$ specified parameters, if the code runs successfully, we use a Python script to assess the correctness of the $k$ specified parameters, resulting in a score of $k/K$. If the code fails to run, the score is zero.
Below is a template for constructing a problem.
In graph theory, several fundamental tasks are crucial for various applications.
Shortest Path algorithms, like Dijkstra’s and Bellman-Ford, help find the minimum distance between nodes in a weighted graph, playing a key role in routing and network design.
Hamiltonian Cycle detection determines whether a cycle exists that visits each vertex exactly once, a problem common in scheduling and logistics.
Subgraph Matching focuses on finding a smaller graph pattern within a larger graph, widely used in chemical informatics and pattern recognition.
These tasks, along with Graph Coloring, Spanning Trees, and Maximum Flow, demonstrate the diversity and utility of analyzing graph structures.
With the rise of Graph Neural Networks (GNNs), new ways of addressing graph-based tasks, like node classification, link prediction, and clustering, have emerged, allowing models to learn from the inherent relationships in graph-structured data.
Task: Given the dataset name dataset, can you use Python library to train a layer number-layer GNN model name model with hidden dimension as hidden dimension, dropout as dropout and predict the accuracy ?
A Python code example for solving GNN-AutoGL problems.
import pkg_resources
import autogl
import torch
from autogl.module.train import Acc
from autogl.solver.utils import get_graph_labels,
get_graph_masks
from autogl.datasets import build_dataset_from_name
cora_dataset = build_dataset_from_name(’cora’)
encoder_hp = {
    ’num_layers’: 2,
    ’hidden’: [128, 256],
    ’dropout’: 0.6,
    ’act’: ’relu’,
    ’eps’: ’false’
predicted = solver.predict_proba()
label = get_graph_labels(cora_dataset[0])
[get_graph_masks(cora_dataset[0], ’test’)].cpu().
numpy()
print(’Test accuracy: ’, Acc.evaluate(predicted,
label))
Appendix C Dataset Statistics
See Table 3.
Appendix D Experimental Results
D.1. NLGraph
See Table 4.
D.2. GraphWiz
See Table 5.
D.3. LLM4DyG
See Table 6.
D.4. GraphInstruct
See Table 7 and Table 8.
D.5. Talk Like a Graph
Coming soon.

### v3

@@
-GraphTeam: Facilitating Large Language Model-based Graph Analysis via Multi-Agent Collaboration
-Abstract.
+GraphTeam: Facilitating Large Language Model-based Graph Reasoning via Multi-Agent Collaboration
+Abstract
 Graphs are widely used for modeling relational data in real-world scenarios, such as social networks and urban computing.
@@
 Extensive experiments on six graph analysis benchmarks demonstrate that GraphTeam achieves state-of-the-art performance with an average 25.85% improvement over the best baseline in terms of accuracy.
-The code and data are available at https://github.com/BUPT-GAMMA/GraphTeam.
-Large Language Model, Multi-agent System, Graph Analysis
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
-1. Introduction
-Graph is a commonly used data structure for modeling relational data.
-Correspondingly, tasks and methods of graph analysis have been extensively exploited in AI field (Brown et al., 2020; Team et al., 2024; Anthropic, 2024) and achieved significantly success in many fields, such as social network (Myers et al., 2014; Min et al., 2021) and unban computing (Yan et al., 2024; Zou et al., 2024).
-Recently, with the surge of LLM, some researchers begin to explore the feasibility of graph analysis with LLMs (Tang et al., 2024; Zhang et al., 2024a; Ouyang et al., 2024; Sanford et al., 2024).
-Existing LLM-based graph analysis methods mainly fall into two categories.
-The first category (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023) usually combines LLMs with graph neural networks (GNNs) for modeling.
-A typical paradigm is to first encode the input graph with a GNN, and then feed the node encodings to an LLM as language token embeddings.
-However, this line of work is specialized for graph machine learning tasks, especially node classification, and has limited transferability to general graph analysis problems.
-The second category (Fatemi et al., 2023; Chen et al., 2024b) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c).
-Strategies like in-context learning (Brown et al., 2020) and chain-of-thought reasoning (Wei et al., 2023) are widely used in these methods to improve performance.
+1 Introduction
+Figure 1: The overall pipeline of our multi-agent system GraphTeam (left), and the comparison between GraphTeam and the best baseline on six benchmarks (right).
+Social networks Garton et al.
+(1997), urban computing Zheng et al.
+(2014), and various other fields make extensive use of graph data and related analytical methods.
+Given the promising problem-solving abilities demonstrated by large language models (LLMs) Brown et al.
+(2020); Anil et al.
+(2024); Anthropic (2024) across different domains, researchers are now exploring the feasibility of using LLMs to reason over graphs Tang et al.
+(2024); Sanford et al.
+(2024); Ouyang et al.
+(2024); Zhang et al.
+(2024a).
+Existing LLM-based graph reasoning methods mainly fall into two categories.
+The first category Chai et al.
+(2023); Liu et al.
+(2023); Qin et al.
+(2023); Liu et al.
+(2024) usually combines LLMs with graph neural networks (GNNs) for modeling.
+A typical paradigm is to first encode the input graph with a GNN, and then feed the encodings to an LLM as language token embeddings.
+However, this line of work is specialized for graph machine learning tasks, especially node classification, and thus has limited transferability.
+The second category Fatemi et al.
+(2024); Chen et al.
+(2024c) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing Liu and Wu (2025); Huang et al.
+(2023); Zhao et al.
+(2024); Shi et al.
+(2023); Chen et al.
+(2024c).
+Strategies like in-context learning Brown et al.
+(2020) and chain-of-thought reasoning Wei et al.
+(2022) are widely used in these methods to improve performance.
 However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
-Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents (Xi et al., 2023; Wang et al., 2024b).
-Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems (Chen et al., 2023; Qian et al., 2024; Hong et al., 2023; Wu et al., 2023), where agents with different specialities can discuss and collaborate with each other to address complex problems.
-However, previous multi-agent frameworks are designed for general purposes and have few knowledge relevant to graph analysis, and thus fail to outperform the aforementioned methods tailored for graph analysis.
-Figure 1. The overall pipeline of our multi-agent system GraphTeam (left), and the comparison between GraphTeam and state-of-the-art baseline on six benchmarks (right).
-To overcome the limitations of previous work, we aim to design a multi-agent system based on LLMs with its design and implementation specialized for graph analysis.
-Inspired by the problem-solving process of humans (Robson and Pólya, 1945; Ramnarine-Rieks, 2011; Bybee et al., 2006; Brand-Gruwel et al., 2009; Savery and Duffy, 1995; Eisenberg and Berkowitz, 1990), we outline our system GraphTeam with three core modules: an input-output normalization module, an external knowledge retrieval module, and a problem-solving module.
-The overall pipeline of GraphTeam is shown in Fig.
-1.
-(1) During human problem-solving processes, clearly understanding and formalizing the problem is the first step (Bybee et al., 2006; Robson and Pólya, 1945), and correctly answering the question in required formats is also important.
-Therefore, we introduce the input-output normalization module with two different LLM-based agents: the question agent aims to extract and refine four key arguments (e.g., graph type and output format) from the original question description, facilitating the system to better understand the problem; the answer agent aims to reformat the results to meet the output requirement.
-(2) During the solving process of a hard problem, humans will seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to previously solved ones (Bybee et al., 2006).
-Therefore, we first build a knowledge base consisting of relevant documentation and experience information.
-Then we introduce the external knowledge retrieval module with a search agent to retrieve the most relevant entries from the knowledge base for each question.
-The retrieved information is expected to benefit the downstream solving process.
-(3) In the problem-solving module, we introduce a coding agent that uses established algorithms via programming to generate solutions.
-If the coding agent fails to generate effective codes, a reasoning agent will directly infer the results without programming.
-We conduct extensive experiments on six graph analysis benchmarks, including Talk like a Graph (Fatemi et al., 2023), LLM4DyG (Zhang et al., 2024b), GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), Graph Instruct (Luo et al., 2024) and a newly introduced GNN-AutoGL based on AutoGL (Guan et al., 2021).
+Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents Xi et al.
+(2023); Wang et al.
+(2024b).
+Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems Qian et al.
+(2024); Hong et al.
+(2024); Wu et al.
+(2024); Chen et al.
+(2024b), where agents with different specialities can collaborate with each other to address complex problems.
+But previous multi-agent frameworks are designed for general purposes, and have few knowledge relevant to graph reasoning.
+Inspired by human problem-solving processes Robson and Pólya (1945); Eisenberg and Berkowitz (1990); Bybee et al.
+(2006); Ramnarine-Rieks (2011), we present GraphTeam, a multi-agent system with three core modules specialized for graph reasoning: (1) Input-output normalization module.
+In human problem solving, clearly understanding and formalizing the problem is the first step Robson and Pólya (1945); Bybee et al.
+(2006), and answering the question in required formats is also important.
+Therefore, we introduce two different LLM-based agents: the question agent extracts and refines key arguments from the original question to help the system better understand the problem; the answer agent reformats the results to meet the specified output requirement.
+(2) External knowledge retrieval module.
+In tackling complex problems, humans often seek help from external sources Eisenberg and Berkowitz (1990) or draw analogies to previously solved ones Bybee et al.
+(2006).
+Therefore, we first build a knowledge base that includes relevant documentation and experience information.
+Then we introduce a search agent to retrieve the most relevant entries from this knowledge base to assist the downstream solving process.
+(3) Problem-solving module.
+Given the retrieved information, we introduce a coding agent that uses established algorithms via programming to generate solutions.
+If the coding agent fails, a reasoning agent directly infers the solution without relying on programming.
+The overall pipeline of GraphTeam is shown in Figure 1.
+We evaluate GraphTeam through extensive experiments across six graph reasoning benchmarks.
 These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs.
-The experimental results and ablation studies demonstrate the effectiveness of GraphTeam as well as the necessity of each module.
+The experimental results and ablation studies highlight the effectiveness of GraphTeam as well as the necessity of each module.
 In summary, our contributions are three-fold:
-$\bullet$ Different from previous paradigms that utilize LLMs for graph analysis, we propose to facilitate the analysis via the collaboration of multiple LLM-based agents.
-$\bullet$ Inspired by the problem-solving process of humans, we propose a carefully designed framework GraphTeam with five agents from three functional groups, i.e., input-output normalization, external knowledge retrieval, and problem-solving modules.
-The agents can work in cooperation with a due division of labor.
-$\bullet$ We conduct extensive experiments over very recent graph analysis benchmarks.
-Experiments show that GraphTeam achieves state-of-the-art (SOTA) performance on all six benchmarks, with an average 25.85% improvement on accuracy over the best baseline.
-2. Related Work
-2.1. LLM-based Graph Analysis
-2.1.1. Benchmarks
-To evaluate the ability of large language models (LLMs) in solving graph analysis problems, researchers have developed various benchmarks (Das et al., 2024; Chen et al., 2024a; Guo et al., 2023; Fatemi et al., 2023).
-Among them, NLGraph (Wang et al., 2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations.
-Graphwiz (Chen et al., 2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete problems.
-GraphInstruct (Luo et al., 2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems.
-Talk like a Graph (Fatemi et al., 2023) measures basic graph understanding abilities, while LLM4DyG (Zhang et al., 2024b) focuses on dynamic graph tasks related to temporal and spatial information.
-In our experiments, we employ the above five benchmarks and also build an additional one focusing on GNNs for assessing the abilities of different methods.
-2.1.2. Methods
-There is a rising trend of research to utilize LLMs for various graph analysis tasks (Li et al., 2024b; Jin et al., 2024).
-Here we focus on the most relevant methods that employ LLMs as predictors (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023).
+$\bullet$ Different from previous paradigms that utilize LLMs for graph reasoning, we propose to facilitate the solving process via the collaboration of multiple LLM-based agents.
+$\bullet$ Inspired by human problem-solving processes, we propose GraphTeam, a carefully designed framework consisting of five agents across three functional groups: input-output normalization, external knowledge retrieval, and problem-solving modules.
+The agents collaborate with a well-defined division of labor for graph reasoning.
+$\bullet$ Extensive experiments show that GraphTeam achieves state-of-the-art (SOTA) performance across all six graph reasoning benchmarks, with an average accuracy improvement of 15.3% over the best baseline.
+2 Related Work
+2.1 LLM-based Graph Reasoning
+To evaluate the ability of LLMs in solving graph reasoning problems, researchers have developed various benchmarks Guo et al.
+(2023); Fatemi et al.
+(2024); Das et al.
+(2024); Chen et al.
+(2024a); Li et al.
+(2024b).
+Among them, NLGraph Wang et al.
+(2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations.
+Graphwiz Chen et al.
+(2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete ones.
+GraphInstruct Luo et al.
+(2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems.
+Talk like a Graph Fatemi et al.
+(2024) measures basic graph understanding abilities, while LLM4DyG Zhang et al.
+(2024b) focuses on dynamic graph tasks related to temporal and spatial information.
+In our experiments, we employ the above five benchmarks, and also construct an additional one to evaluate the ability of deploying GNNs with AutoGL Guan et al.
+(2021).
+There is a rising trend of research to utilize LLMs for various graph reasoning tasks Li et al.
+(2024c); Jin et al.
+(2024).
+Here we focus on the most relevant methods that employ LLMs as predictors Chai et al.
+(2023); Liu et al.
+(2023); Qin et al.
+(2023); Liu et al.
+(2024).
 Related work can be roughly divided into two categories.
-The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification (Tang et al., 2024; Zhang et al., 2024a).
-For example, GraphGPT (Tang et al., 2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification.
-GraphTranslator (Zhang et al., 2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification.
-However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse analysis tasks in the above benchmarks.
-The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning (Brown et al., 2020), chain-of-thought reasoning (Wei et al., 2023) for inference (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c).
-For example, GUNDAM (Ouyang et al., 2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference.
-However, recent studies have shown that the reasoning depths of current LLMs are still shallow (Lam et al., 2024; Li et al., 2024a), and will probably fail for graph analysis problems that require complex reasoning.
+The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification Tang et al.
+(2024); Zhang et al.
+(2024a).
+For example, GraphGPT Tang et al.
+(2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification.
+GraphTranslator Zhang et al.
+(2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification.
+However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse reasoning tasks in the above benchmarks.
+The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning Brown et al.
+(2020), chain-of-thought reasoning Wei et al.
+(2022) for inference Liu and Wu (2025); Huang et al.
+(2023); Zhao et al.
+(2024); Shi et al.
+(2023); Chen et al.
+(2024c).
+For example, GUNDAM Ouyang et al.
+(2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference.
+However, recent studies have shown that the reasoning depths of current LLMs are still shallow Parmar et al.
+(2024); Li et al.
+(2024a), and will probably fail for problems that require complex reasoning.
 In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
-2.2. Multi-Agent Collaboration of LLMs
+2.2 Multi-Agent Collaboration of LLMs
 Autonomous agents aim to solve problems with self-directed planning and actions.
-Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions (Xi et al., 2023; Bang et al., 2023; Li et al., 2023; Wang et al., 2024b).
-Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where multiple agents work in cooperation with a due division of labor.
-For example, ChatDev (Qian et al., 2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration.
-MetaGPT (Hong et al., 2023) is a meta-programming framework that simulates human workflows.
+Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions Bang et al.
+(2023); Li et al.
+(2023); Xi et al.
+(2023); Wang et al.
+(2024b).
+Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where agents collaborate with a well-defined division of labor.
+For example, ChatDev Qian et al.
+(2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration.
+MetaGPT Hong et al.
+(2024) is a meta-programming framework that simulates human workflows.
 With proper developing pipeline, MetaGPT can generate effective programming solutions.
-AgentVerse (Chen et al., 2023) is a multi-agent framework inspired by human collaboration dynamics.
+AgentVerse Chen et al.
+(2024b) is a multi-agent framework inspired by human collaboration dynamics.
 It involves four iterative stages of expert recruitment, decision discussion, action and evaluation.
-AutoGen (Wu et al., 2023) is another multi-agent framework for general purposes.
+AutoGen Wu et al.
+(2024) is another multi-agent framework for general purposes.
 Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks.
-MACRec (Wang et al., 2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration.
-This system includes various specialized agents and can easily solve various recommendation tasks.
-However, these multi-agent frameworks are not specialized for graph analysis, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems.
-Consequently, they can hardly beat the SOTA performance on the aforementioned graph analysis benchmarks.
-3. Methodology
-In this section, we propose GraphTeam, a multi-agent system based on LLMs for graph analysis.
-We will start by introducing the overall framework, and then present the details of each agent in the system.
-3.1. Framework Overview
-As shown in Fig.
-2, the entire framework of GraphTeam consists of five agents from three functional groups, i.e., question and answer agents for input-output normalization, search agent for external knowledge retrieval, as well as coding and reasoning agents for problem solving.
-Figure 2. The overall framework of GraphTeam, which includes five agents from three functional groups.
+MACRec Wang et al.
+(2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration.
+However, these multi-agent frameworks are not specialized for graph reasoning, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems.
+3 Methodology
+We start by presenting the overall framework of GraphTeam, and then present the details of each agent in the system.
+3.1 Framework Overview
+As shown in Figure 2, the entire framework of GraphTeam consists of five agents across three functional groups:
+Figure 2: The overall framework of GraphTeam, which includes five agents across three functional groups.
+Firstly, the question agent refines the original question by extracting key arguments.
+Then the search agent queries knowledge bases with the reformatted question to retrieve relevant experience and documentation.
+Subsequently, the coding agent attempts to solve the problem via programming.
+If the coding agent fails after several retries, the reasoning agent will take over and directly solve the question without programming.
+Finally, the solutions from either the coding or reasoning agents are then passed to the answer agent for format standardization.
 Input-Output Normalization: This functional group aims to refine key information from input questions, and map the results to desired output formats.
@@
 External Knowledge Retrieval: This functional group only includes the search agent to extract relevant information from external knowledge base.
-Specifically, we first build a knowledge base for graph analysis based on documentation of popular Python libraries and experiences of previous problem-solving processes.
+Specifically, we first build a knowledge base for graph reasoning based on documentation of popular Python libraries and experiences of previous problem-solving processes.
 With problems refined by the upstream question agent as queries, the search agent then retrieves information beneficial to downstream problem solving from the knowledge base.
-Problem Solving: This functional group aims to calculate the results based on the refined problem of question agent and the extracted knowledge of search agent.
-Specifically, the coding agent attempts to write Python codes to solve the problem.
-When the generated codes cannot be executed normally, a retry mechanism will be employed to fix the errors in the codes.
-If the coding agent still fails after several retries, the reasoning agent will solve the problem directly without programming.
-3.2. Question Agent
-Abstraction is a preliminary step for human beings to solve problems (Robson and Pólya, 1945).
+Problem Solving: This functional group calculates the results based on the refined question of question agent and the extracted knowledge of search agent.
+Specifically, the coding agent attempts to solve the problem via programming.
+If the coding agent fails after several retries, the reasoning agent will solve the problem directly without programming.
+3.2 Question Agent
+Abstraction is a preliminary step for human beings to solve problems Robson and Pólya (1945).
 Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions.
-Formally, given a graph analysis problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
-(1) $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q),$
-where $q_{r}$ is a condensed version of original instruction, $q_{t}$ indicates the graph type (e.g., directed, undirected, weighted, etc.), $q_{g}$ contains the graph to be analyzed, and $q_{f}$ specifies how the answer should be presented (e.g., numerical value, list, etc.).
-By explicitly identifying these key information, downstream agents can better understand a task and its requirements, thereby enhancing the overall analysis capability.
-An example of the question agent is shown below:
+Formally, given a graph reasoning problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
+$(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q),$ (1)
+where $q_{r}$ is a condensed version of original instruction, $q_{t}$ indicates the graph type (e.g., directed, undirected, weighted, etc.), $q_{g}$ contains the graph to be processed, and $q_{f}$ specifies how the answer should be presented (e.g., numerical value, list, etc.).
+By explicitly identifying these key information, downstream agents can better understand a task and its requirements, thereby enhancing the overall reasoning capability.
+An example of question agent is shown below:
 Original Question
-In an undirected graph … The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5).
-Q: Is there a path in this graph that visits every node exactly once? …
-A: In the answer, …, Output format such as: … 1, 4, 8.
+In an undirected graph …, the edges are: (0,3) … (4,5).
+Q: Is there a path … visits every node exactly once? …
+A: In the answer, …, Output format such as : …
 Reformatted Question
-refined_question: Is there a path in this graph that visits …
+refined_question: Is there a path in this graph that …
 graph_type: undirected
-input_data: ’Nodes’: [0, 1, 2, 3, 4, 5, 6],
-’Edges’: [[0, 3], [0, 1], [1, 4], [1, 6], [2, 4], [3, 5], [3, 6], [4, 5]]
+input_data: Nodes: [0, 1, 2, 3, 4, 5, 6],
+Edges: [0, 3], [0, 1], [1, 6], [2, 4], [3, 5], [3, 6], [4, 5]
 output_format: Yes. The path can be: 1,4,8.
-3.3. Search Agent
-During the solving process of a hard problem, it is common for human beings to seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to a previously solved one (Bybee et al., 2006).
-Therefore, we first build a knowledge base with documentation and experience information to enhance the programming ability of the multi-agent system, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
-3.3.1. Knowledge Base Construction
+3.3 Search Agent
+During the solving process of a hard problem, it is common for human beings to seek help from external sources Eisenberg and Berkowitz (1990) or make analogy to a previously solved one Bybee et al.
+(2006).
+Therefore, we first build a knowledge base with documentation and experience information, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
+3.3.1 Knowledge Base Construction
 The knowledge base has two parts, i.e., documentation base $\mathcal{K}_{doc}$ and experience base $\mathcal{K}_{exp}$. Examples of documentation and experience entries are shown below:
@@
 api_name: hamiltonian_path
-description: Returns a Hamiltonian path in the given …
-parameters: ”G : NetworkX graph” , …
-returns: ”path : list”: ”A list of nodes which form a Hamiltonian path in G.”
-examples:
-    import NetworkX as nx
-    nx.tournament.hamiltonian_path(G)
+description: Returns a Hamiltonian path …
+parameters: G : NetworkX graph, …
+returns: path: list…
+examples: …
+Experience Information
+question: In an undirected graph, …, the edges are: (0,1) … (5,6) Q: Is there a path in this graph …
+answer: Yes. The path can be: 0, 1, 6, 5.
+thought: This is finding a Hamiltonian path. …
+code:
+    def hamiltonian_path(graph, path):
 Documentation.
@@
 Only the experience with the highest utility will be kept in each $\mathcal{K}_{exp}^{t}$.
-[…]
-Experience Information
-question: In an undirected graph, …, The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5) Q: Is there a path in this graph …
-answer: Yes. The path can be: 0, 1, 6, 3, 5, 4, 2.
-meta information:
-thought: The experience from this task involved solving the problem of finding a Hamiltonian path. …
-code:
-    import NetworkX as nx
-    def hamiltonian_path_backtrack(graph, path):
-3.3.2. Knowledge Retrieval
-For effective information retrieval, we build vector databases with LlamaIndex (lla, 2024) for documentation and experiences, respectively.
+More details including the pseudo codes for building the experience base can be found at Alg.
+2 in Appendix A.1.
+3.3.2 Knowledge Retrieval
+For effective information retrieval, we build vector databases with LlamaIndex Liu (2022) for documentation and experiences, respectively.
 Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database.
@@
 For simplicity, we denote the operation of the search agent as
-(2) $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta).$
-Algorithm 1 Problem-Solving Experience Collection
-1:Training dataset $\mathcal{D}_{\text{train}}$, validation dataset $\mathcal{D}_{\text{valid}}$, candidate experience size $N_{exp}$ for each problem type $t\in\mathcal{T}$.
-2:Experience knowledge base $\mathcal{K}_{exp}$.
-3:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{train}}$ do
-4:     if $|\mathcal{K}_{exp}^{t}|$ $N_{exp}$ then
-[…]
-6:         if answer equals to ground truth $a$ then
-[…]
-8:         end if
-9:     end if
-10:end for
-11:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{valid}}$ do
-12:     for $e\in\mathcal{K}_{exp}^{t}$ do
-[…]
-14:         if answer equals to ground truth $a$ then
-15:              $\text{Utility}_{e}\leftarrow\text{Utility}_{e}+1$
-16:         end if
-17:     end for
-18:end for
-19:for $t\in\mathcal{T}$ do
-20:     $\mathcal{K}_{exp}^{t}\leftarrow\text{argmax}_{e\in\mathcal{K}_{exp}^{t}}\text{Utility}_{e}$
-21:end for
-22:$\mathcal{K}_{exp}\leftarrow\{\mathcal{K}_{exp}^{t}\}_{t\in\mathcal{T}}$
-3.4. Coding and Reasoning Agents
+$\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta).$ (2)
+3.4 Coding and Reasoning Agents
 Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems.
 In practice, the generated codes may encounter compilation or running errors.
-Following the trial-and-error strategy (Robson and Pólya, 1945; Ramnarine-Rieks, 2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages.
+Following the trial-and-error strategy Robson and Pólya (1945); Ramnarine-Rieks (2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages.
 More formally, the operation of coding agent in the $n$-th trial can be written as:
-(3) $\begin{split}(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},\text{knowledge}_{q},\\ \{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1}),\end{split}$
+$\begin{split}(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},\\ \text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1}),\end{split}$ (3)
 where $\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n}$ are respectively the execution result, generated code, and error message of the $n$-th trial.
@@
 If the codes runs normally, we define $\text{error}_{q}^{n}$ as $\tt None$.
-If the coding agent fails to fix the codes before reaching the maximum number of trials, we will employ the reasoning agent to directly answer the question without programming:
-(4) $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f}).$
-3.5. Answer Agent
+If the coding agent fails to fix the codes when reaching the maximum number of trials, we will employ the reasoning agent to directly answer the question without programming:
+$\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f}).$ (4)
+In this way, the coding and reasoning agents can solve problems collectively, and achieve a collaborative functionality at the system level.
+3.5 Answer Agent
 Based on the output format $q_{f}$, the answer agent will organize the results given by the coding or reasoning agents as final results.
 Similar to the retry mechanism of coding agent, we also introduce a self-checking mechanism that iteratively refines the results according to format requirement $q_{f}$.
-Formally, the operation of answer agent can be written as:
-(5) $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1}),$
+Formally, the operation of answer agent is:
+$\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1}),$ (5)
 where $\text{output}_{q}^{n}$ is the output in the $n$-th iteration, and $\text{output}_{q}^{0}$ is $\text{result}_{q}$ generated by the coding or reasoning agents.
@@
 result: Yes. The path can be: 0, 1, 6, 3, 5, 4, 2.
-3.6. Summary
+3.6 Summary
 We formalize the inference pipeline of GraphTeam in Alg.
-2.
+1.
 Most agent functions (including Question, Coding, Reasoning and Answer) in Alg.
 1 are all implemented by an LLM with different prompt templates.
-The Search function is implemented based on LlamaIndex (lla, 2024), and the Solver function in Alg.
-1 is implemented by a simplified version of GraphTeam without experience base.
-We will provide more implementation details in the appendix.
+The Search function is implemented based on LlamaIndex Liu (2022).
+We will provide more implementation details in Appendix A.
 Algorithm 1 Inference Pipeline of GraphTeam
-1:Problem described by question $q$, knowledge base $\mathcal{K}$, threshold $\delta$ for determining whether experience or documentation is retrieved, maximum number of retries $N_{retry}$, number of self-checking iterations $N_{check}$.
-2:Final answer of the problem $\text{output}_{q}$.
-3:$(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
-4:$\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
-5:$n=0$
-6:while $\text{error}_{q}^{n}\neq\tt{None}$ and $n<N_{retry}$ do
-7:     ++n
-8:     $(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},$ $\text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1})$
-9:     $\text{result}_{q}\leftarrow\text{result}_{q}^{n}$
-10:end while
-11:if $\text{error}_{q}^{n}\neq\tt{None}$ then
-12:     $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f})$
-13:end if
-14:$\text{output}_{q}^{0}\leftarrow\text{result}_{q}$
-15:for $n=1,2,\dots,N_{check}$ do
-16:     $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1})$
-17:end for
-4. Experiments
-In this section, we conduct experimental evaluation to answer the following research questions (RQs): RQ1: How effective is our proposed GraphTeam compared to the state-of-the-art (SOTA) baselines?
-RQ2: Have each component of GraphTeam played their roles effectively?
-RQ3: How does our framework perform with respect to different task categories and output formats of graph analysis problems?
-RQ4: How do the hyper-parameters of certain components in the system have an impact on the overall performance?
-4.1. Experimental Setup
-4.1.1. Datasets
-To evaluate the effectiveness of our system in solving graph analysis problems, we use six benchmarks for LLM-based graph analysis: Talk like a Graph (Fatemi et al., 2023) primarily studies some fundamental problems in graph analysis and explores the cognitive abilities of LLMs regarding graphs.
-LLM4DyG (Zhang et al., 2024b) focuses on basic problems related to dynamic graphs, while GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), and Graph Instruct (Luo et al., 2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory.
-GNN-AutoGL (details shown in Appendix B) is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL (Guan et al., 2021).
-4.1.2. Metrics
-For benchmarks with definite answers (Fatemi et al., 2023; Chen et al., 2024a; Wang et al., 2024a; Luo et al., 2024; Zhang et al., 2024b), we use accuracy as the evaluation metric.
+0:  Problem described by question $q$, knowledge base $\mathcal{K}$, threshold $\delta$ for determining whether experience or documentation is retrieved, maximum number of retries $N_{retry}$, number of self-checking iterations $N_{check}$.
+0:  Final answer of the problem $\text{output}_{q}$.
+1:  $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
+2:  $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
+3:  $n=0$
+4:  while $\text{error}_{q}^{n}\neq\tt{None}$ and $n<N_{retry}$ do
+5:     ++n
+6:     $(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},$ $\text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1})$
+7:     $\text{result}_{q}\leftarrow\text{result}_{q}^{n}$
+8:  end while
+9:  if $\text{error}_{q}^{n}\neq\tt{None}$ then
+10:     $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f})$
+11:  end if
+12:  $\text{output}_{q}^{0}\leftarrow\text{result}_{q}$
+13:  for $n=1,2,\dots,N_{check}$ do
+14:     $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1})$
+15:  end for
+4 Experiments
+We conduct extensive experiments on six graph reasoning benchmarks to validate the effectiveness of GraphTeam.
+4.1 Experimental Setup
+4.1.1 Datasets
+To evaluate the effectiveness of our system, we consider six benchmarks for LLM-based graph reasoning: Talk like a Graph Fatemi et al.
+(2024) primarily studies some fundamental problems in graph reasoning and explores the cognitive abilities of LLMs regarding graphs.
+LLM4DyG Zhang et al.
+(2024b) focuses on basic problems related to dynamic graphs, while GraphWiz Chen et al.
+(2024a), NLGraph Wang et al.
+(2024a), and GraphInstruct Luo et al.
+(2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory.
+GNN-AutoGL is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL Guan et al.
+(2021).
+Dataset statistics and details can be found at Appendix B.1 and  B.2.
+4.1.2 Metrics
+For benchmarks with definite answers Fatemi et al.
+(2024); Wang et al.
+(2024a); Zhang et al.
+(2024b); Chen et al.
+(2024a); Luo et al.
+(2024), we use accuracy as the evaluation metric.
 For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria.
-Specifically, we extract the generated code from the model response and compare the execution result with the reference solution to assess whether the model-generated results match the given standard answers or reference solutions.
-For the first five benchmarks (Fatemi et al., 2023; Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024; Wang et al., 2024a), the evaluation results are binary based on the exact match between the execution results and the standard answers or reference solutions.
-For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0.
-Otherwise, we check whether the hyper-parameters(eg.
-model name, dataset name, hidden dimension, dropout) meet the requirements.
-Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
-4.1.3. Settings
-We choose GPT-4o-mini (OpenAI, 2024) as the base LLM for all agents of our GraphTeam.
-For Python libraries, we use NetworkX version 3.3111https://networkx.org/ and AutoGL version 0.4.0222https://mn.cs.tsinghua.edu.cn/autogl/documentation/, together with Python version 3.10.14.
+More details are shown in Appendix B.3.
+4.1.3 Settings
+We choose GPT-4o-mini Hurst et al.
+(2024) and Qwen2.5-Coder-7B Hui et al.
+(2024) as the base LLM for the agents of GraphTeam.
+For Python libraries, we use NetworkX version 3.3 Hagberg et al.
+(2008) and AutoGL version 0.4.0 Guan et al.
+(2021), together with Python version 3.10.14.
 The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$.
 The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
-4.1.4. Baselines
-As the SOTA methods for each benchmark are quite diverse and some of which are specialized for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM (Ouyang et al., 2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers (Sanford et al., 2024) as the SOTA; for GraphWiz, LLM4DyG, and GraphInstruct (Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024), we select the best model in their benchmark papers.
-For GraphWiz (Chen et al., 2024a), we did not choose GraphAgent-Reasoner (Hu et al., 2024) because it only provide results for 6 out of 9 categories.
-4.2. Main Results (RQ1)
-Table 1. Performance comparison on six graph analysis benchmarks in terms of accuracy (%).
-Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
-Main GPT-4o-mini 71.67 47.61 51.40 58.03 53.67 3.28
-SOTA 97.30 65.00 79.70 53.71 43.53 -
-GraphTeam 99.13 88.61 97.80 94.81 88.14 48.44
-     - question agent 85.87 62.81 79.22 59.18 85.86 14.53
-Input-Output Normalization      - answer agent 94.80 87.97 96.70 94.27 85.25 45.78
-     - documentation 92.03 87.97 96.40 94.07 86.12 46.88
-Knowledge Retrieval      - experience 89.96 82.83 54.70 60.26 76.64 4.76
-     - coding agent 69.40 46.39 39.30 59.33 61.09 -
-     - retry 91.90 83.53 87.70 89.33 85.66 44.22
-Problem Solving      - reasoning agent 92.77 54.17 96.90 94.00 85.45 46.36
-Table 2. Performance comparison on the hard subset of six graph analysis benchmarks in terms of accuracy (%).
-Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
-Main GPT-4o-mini 47.90 28.92 44.93 42.56 23.19 3.28
-SOTA 93.90 70.50 44.96 33.90 11.71 -
-GraphTeam 99.20 76.25 98.14 98.33 73.33 48.44
-     - question agent 70.00 39.75 69.82 53.67 73.29 14.53
-Input-Output Normalization      - answer agent 97.40 75.25 97.51 96.33 70.00 45.78
-     - documentation 98.90 75.25 97.51 94.33 69.76 46.88
-Knowledge Retrieval      - experience 98.40 62.92 42.32 70.56 63.95 4.76
-     - coding agent 64.00 38.33 19.21 51.56 38.24 -
-     - retry 95.40 61.00 40.27 85.44 70.95 44.22
-Problem Solving      - reasoning agent 98.90 66.08 97.51 95.67 69.62 46.36
-We present the results in Tab.1.
-GraphTeam surpasses previous SOTA by 1.83-44.61% in accuracy, and outperforms the base LLM (GPT-4o-mini) by 27.46-46.40% in accuracy.
-Such significant improvements fully demonstrate the effectiveness of our carefully designed multi-agent system.
-Besides, the time cost for answering each problem is less than 3 seconds, and the monetary cost of API usage is less than \$0.01.
-To further demonstrate GraphTeam’s performance on challenging problems, we assess its performance on the “hard” subset of each benchmark.
-For GraphWiz, we adhere to the official partition and report results on their designated hard set.
-For other benchmarks, which all have their problems categorized into different types or domains, we employ a model-based approach: we identify the three categories where previous SOTA performs poorest and select problems from these categories.
-This selection method ensures we target the most challenging aspects of each benchmark.
-Tab.
-2 summarizes the system’s performance on these high-difficulty questions.
-Notably, GraphTeam maintains high accuracy on the Talk like a Graph hard set (99.20%), and shows significant improvements over SOTA for NLGraph (98.14% v.s.
-44.96%) and LLM4DyG (98.33% v.s.
-33.90%).
-While performance drops are observed for GraphWiz and GraphInstruct compared to the general set, GraphTeam still substantially outperforms both GPT-4o-mini and SOTA methods.
-This consistent and strong performance across diverse hard sets underscores GraphTeam’s robustness and its ability to effectively handle complex graph-related tasks.
-4.3. Ablation Study (RQ2)
-To evaluate the contribution of each component, we conduct a systematic ablation study.
+4.1.4 Baselines
+As the SOTA methods for each benchmark are quite diverse and some of which are designed for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM Ouyang et al.
+(2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers Sanford et al.
+(2024) as the SOTA; for LLM4DyG, GraphWiz, and GraphInstruct Zhang et al.
+(2024b); Chen et al.
+(2024a); Luo et al.
+(2024), we select the best model in their benchmark papers.
+In addition, we adopt AutoGPT Significant Gravitas (2023) and Open Interpreter Open Interpreter (2023) as single-agent baselines, while employing AutoGen Wu et al.
+(2024) and ChatDev Qian et al.
+(2024) as multi-agent baselines for comparison.
+4.2 Main Results
+Table 1: Performance comparison between GraphTeam and baselines in terms of accuracy (%).
+Base LLM       Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL Average Rank
+-       Previous SOTA 79.7 90.2 43.5 65.0 53.7 - 5.5
+      Base Model 51.4 61.3 53.7 47.6 58.0 3.3 7.5
+      AutoGPT 65.8 65.2 58.5 50.2 61.6 10.3 5.7
+      Open Interpreter 68.2 85.6 54.2 53.9 55.5 8.8 5.5
+      AutoGen 59.7 70.2 57.7 62.8 47.3 28.1 5.8
+      ChatDev 79.6 72.5 70.6 67.8 77.8 34.5 3.5
+GPT-4o-mini       GraphTeam 97.8 94.5 88.1 88.6 94.8 48.4 1.0
+      Base Model 37.3 63.1 34.5 51.8 38.2 1.9 8.3
+Qwen2.5-Coder-7B       GraphTeam 93.0 85.6 84.5 78.6 87.4 36.7 2.2
+As shown in Table 1, GraphTeam achieves significant and consistent improvements over baseline methods on all six benchmarks.
+Compared with the best baseline, GraphTeam (GPT-4o-mini) has 15.3% accuracy gains on average.
+Even when equipped with open-source model, GraphTeam (Qwen2.5-Coder-7B) still has 7.5% average accuracy improvements over the best baseline based on GPTs.
+These results demonstrate that GraphTeam can achieve promising performance with both open-source and closed-source LLMs.
+Among the baselines, LLM-based agents (i.e., AutoGPT, Open Interpreter, AutoGen and ChatDev) consistently achieve better performance than the corresponding base LLM (i.e., GPT-4o-mini).
+For example, the single-agent baseline AutoGPT has 6.1% average accuracy improvements over the base LLM.
+Moreover, the multi-agent baseline ChatDev achieves even better results, with 21.3% average accuracy gains over the base LLM.
+These observations support our motivation of employing LLM-based agents for graph reasoning.
+4.3 Ablation Study
+To evaluate the contribution of each component, we conduct a systematic ablation study of GraphTeam using GPT-4o-mini.
 We assess the impact of Input-Output Normalization by removing the question or answer agent.
@@
 This comprehensive ablation enables us to quantify each component’s significance to the system performance.
-The overall results are presented in Tab.1 and 2.
-Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.44%) by breaking the problem into individual components.
-For a few benchmark such as GraphInstruct (Luo et al., 2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.28%).
-The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.54%- 3.33%).
+The results are presented in Table 2.
+Table 2: Performance comparison between GraphTeam and ablation variants in terms of accuracy (%).
+Category Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
+GraphTeam (GPT-4o-mini) 97.8 94.5 88.1 88.6 94.8 48.4
+     - question agent 79.2 82.1 85.9 62.8 58.6 14.8
+Input-Output Normalization      - answer agent 96.7 93.9 85.3 88.0 94.2 45.8
+     - documentation 96.4 93.8 86.1 88.0 94.0 46.9
+Knowledge Retrieval      - experience 54.7 92.0 76.6 82.8 60.3 4.8
+     - coding agent 39.3 60.7 61.1 46.4 59.3 -
+     - retry 87.7 93.5 85.7 83.5 88.5 44.3
+Problem Solving      - reasoning agent 96.9 93.9 85.5 54.2 93.7 46.4
+Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.3%) by breaking the problem into individual components on average.
+For a few benchmark such as GraphInstruct Luo et al.
+(2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.2%).
+The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.6-2.8%).
 External Knowledge Retrieval: The search agent plays a significant role.
-When solving graph analysis problems, providing the large language model with prior knowledge of related problem-solving experiences or API documentation can help the model solve the problem more effectively.
-The impact of experience is greater than that of documentation, indicating that LLMs solve problems in a way similar to humans with an average improvement of 24.63%.
+When solving graph reasoning problems, providing LLM with related problem-solving experiences or API documentation can help the model solve the problem more effectively.
+The impact of experience is greater than that of documentation, indicating that LLMs solve problems in a way similar to humans with an average improvement of 23.5%.
 They can derive better solutions from prior experiences, effectively addressing similar problems.
-Documentation can enhance the model’s programming capabilities.
-When relevant experiences are not available, documentation is still needed to learn graph analysis methods and subsequently solve corresponding problems (0.64% to 7.1%).
+When relevant experiences are not available, documentation can also help enhance the model’s programming capabilities in solving corresponding problems (0.6% to 2.0%).
 Problem Solving: The coding agent plays a crucial role.
-Without programming and relying solely on reasoning, the system’s performance would significantly decrease, indicating that programming is an effective approach for solving graph analysis problems (27.05% to 58.50%).
-The retry mechanism does not always have a significant overall effect, but it shows clear benefits for hard problems with a maximum improvement of 57.87%.
-This indicates that Hard problems require more complex code, which cannot be written correctly in a single attempt and needs multiple tries to solve effectively.
-The reasoning agent does not have a significant impact on problems where programming achieves good results, but it plays an important role in problems where programming is insufficient and reasoning is required (0.81% to 34.44%).
+Without programming and relying solely on reasoning, the system’s performance would significantly decrease, indicating that programming is an effective approach for solving graph reasoning problems (27.0% to 58.5%).
+The retry mechanism also brings a positive effect (1.0% to 10.1%) by debugging programs.
+The reasoning agent does not have a significant impact on problems where programming achieves good results, but it plays an important role in problems where programming is insufficient and reasoning is required (0.9% to 34.4%).
 The contribution of each component of the system varies across different benchmarks.
-Since the core problem-solving component of this system is the coding agent, programming to solve problems is a crucial and effective approach for graph analysis.
+Since the core problem-solving component of this system is the coding agent, programming to solve problems is a crucial and effective approach for graph reasoning.
 For GNNs, in particular, coding is an indispensable component.
-Therefore, removing the coding agent leads to a significant drop in performance for all benchmarks (42.24% in average).
-Some benchmarks do not have strict requirements for the answer format (NLGraph, GraphWiz, and LLM4DyG), so removing the answer agent does not significantly impact the overall system performance (0.76% in average).
-Since Talk like a Graph contains simple questions, most problems can be solved correctly through coding without the need for the reasoning agent, hence the benchmark has minimal need for the reasoning agent (6.36%).
-Since GNN-AutoGL lacks documentation information, this module does not affect system performance (1.56%).
-Similarly, GraphInstruct’s problem-solving does not rely on documentation, so documentation has little impact on the system’s performance on this benchmark (0.04%).
-Figure 3. Performance with respect to different task categories.
-Figure 4. Performance with respect to different output formats.
-Figure 5. Hyper-parameter analysis of four hyper-parameter in the proposed GraphTeam.
-4.4. Analysis on Different Groupings (RQ3)
-To provide a more nuanced understanding of GraphTeam’s capabilities and limitations, we conduct a detailed analysis of its performance across various task categories and output formats.
-This fine-grained approach allows us to identify specific strengths and weaknesses.
+Therefore, removing the coding agent leads to a significant drop in performance for all benchmarks (40.9% on average).
+Some benchmarks do not have strict requirements for the answer format (Talk like a graph, GraphWiz, and LLM4DyG), thus removing the answer agent does not significantly impact the overall system performance.
+Since questions in Talk like a Graph are relatively simple, most problems can be correctly solved through coding without the need for the reasoning agent.
+Hence this benchmark has minimal need for the reasoning agent.
+In summary, the full model always perform the best, validating the necessity of each design.
+4.4 Analysis on Different Groupings
 Group by Task Categories.
-We propose a categorization scheme that covers a broad spectrum of graph-related problems.
-We classify the questions into four categories based on the nature of the task: (1) basic: fundamental graph understanding, (2) algorithm-macro-level: analysis of entire graph structures, (3) algorithm-micro-level: local properties of nodes and edges, and (4) GNN: graph neural networks.
-As shown in Fig.
-3, GraphTeam achieves an accuracy of over 90% on both basic and macro-level algorithms, and 85.2% on micro-level tasks, indicating that the system can effectively handle common graph analysis tasks.
-However, all systems, including GraphTeam, struggle with GNN-related tasks, suggesting a need for improvement in this area.
-These results reveal interesting insights: For micro-level tasks, removing the reasoning agent causes a significant 21.2% drop in performance, highlighting the complexity of these problems and the limitations of the coding agent alone.
-In GNN-related tasks, while the coding agent and experience are important, the question agent plays a crucial role (38.8% impact when removed).
-This underscores the importance of effectively extracting key information from GNN problem statements.
-These findings point to areas for focused improvement, particularly in enhancing reasoning capabilities for micro-level tasks and information extraction for GNN-related problems.
-Group by Output Formats.
-To account for the diversity in problem outputs, we categorize the questions based on their required response formats: (1) yes/no, (2) digits, (3) list/set, and (4) others.
-This classification allows us to assess system performance across different types of graph analysis outputs.
-As shown in Fig.
-4, GraphTeam demonstrates strong performance on yes/no, digits, and list/set outputs, achieving an accuracy of over 85% for each of these categories.
-However, the system’s performance drops notably for more complex output formats, with an accuracy of only 70.1%.
-This decline in accuracy for the “others” format indicates that GraphTeam, particularly its answer agent, needs enhancement in handling advanced data structures and more complex output formats.
-Improving the system’s capability to generate and structure complex outputs could significantly boost its overall effectiveness in addressing a wider range of graph analysis problems.
-4.5. Hyper-parameter Sensitivity (RQ4)
-We conduct a study on key hyper-parameters using the NLGraph (Wang et al., 2024a) benchmark to determine optimal settings or optimization strategies.
-Fig.
-5 presents our analysis of several important hyper-parameters.
+We classify the questions into four categories based on the nature of the task: (1) basic: fundamental graph understanding, (2) macro-level: reasoning over entire graph structures, (3) micro-level: local properties of nodes and edges, and (4) GNN: graph neural networks.
+As shown in Figure 3, GraphTeam (GPT-4o-mini) achieves over 90% accuracies on the first two categories, and 85.6% on micro-level tasks, ranking the first over all four groups.
+While GraphTeam (Qwen2.5-Coder-7B) always ranks the second, achieving over 76% on the first three categories.
+These results indicate that our framework can effectively handle common graph reasoning tasks.
+For most categories (basic, micro-level, and GNN), we observe consistent improvements when utilizing agent frameworks.
+But on macro-level tasks, we observe that AutoGPT, Open Interpreter, and AutoGen perform worse than using GPT-4o-mini alone.
+We find that this degradation is caused by misunderstanding question requirements during long reasoning processes.
+Thus, employing the question agent to explicitly identify key information is necessary for graph reasoning.
+Besides, all systems struggle with GNN-related tasks, suggesting a need for improvement in future work.
+Figure 3: Performance with respect to different task categories.
+Group by Output Types.
+We categorize problems according to their required data formats: true/false questions (yes/no), numerical calculations (digits), analytical questions (list/set), and other types, with results presented in Figure 4.
+Similarly, GraphTeam with GPT-4o-mini and Qwen2.5-Coder-7B always rank the first and second across all four output formats.
+In particular, GraphTeam (GPT-4o-mini) reaches or approaches 90.0% for the first three categories, showing a strong generalization ability.
+The base LLMs perform good for true/false questions, but are inferior to agent-based methods for problems requiring more complex output formats.
+Figure 4: Performance with respect to different output formats.
+Figure 5: Average accuracies (%) and token consumptions of different agent-based methods.
+While maintaining comparable token consumption with typical agent frameworks, GraphTeam substantially improves the average accuracy for graph reasoning.
+4.5 Efficiency Analysis
+To verify whether our proposed GraphTeam is cost effective, we present average accuracy and token consumption for all agent-based methods based on GPT-4o-mini in Figure 5.
+GraphTeam achieves an average accuracy of 85.4% with relatively low token consumption.
+The token cost of GraphTeam is competitive with single-agent baselines, and is significantly lower than ChatDev.
+Though ChatDev demonstrates notable accuracy advantages over other agent frameworks, it consumes approximately 5.6 times more tokens than GraphTeam.
+This indicates that our framework design is better optimized for graph reasoning tasks, achieving promising performance with acceptable resource consumption.
+4.6 Hyper-parameter Analysis
+As shown in Figure 6, we conduct experiments to study the influence of key hyper-parameters using the NLGraph Wang et al.
+(2024a) benchmark.
 For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance.
@@
 These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
-5. Conclusion
-To address the limitations of previous LLM-based graph analysis methods, we draw inspiration from human problem-solving processes, and design a multi-agent system named GraphTeam based on LLMs.
-GraphTeam consists of five agents from three key modules, namely input-output normalization, external knowledge retrieval, and problem solving.
-The agents can cooperate with each other for solving complex graph analysis problems.
-Extensive experiments on six benchmarks demonstrate that GraphTeam achieves state-of-the-art results with an average improvement of 25.85% in terms of accuracy.
-For future work, we will consider to fine-tune the LLMs in GraphTeam for better performance, and further optimize the system pipeline to accommodate more tasks.
-Appendix A Prompt Templates
-A.1. Question Agent
-Requirement Analyst
-Here is the task: task.
-As the assistant_role, you need to extract key information from the task, specifically focusing on the following elements:
+Figure 6: Analysis of key hyper-parameters in the proposed GraphTeam.
+5 Conclusion
+In this work, we draw inspiration from human problem-solving processes, and design a multi-agent system named GraphTeam for graph reasoning.
+GraphTeam consists of five agents across three key modules, namely input-output normalization, external knowledge retrieval, and problem solving.
+The agents can cooperate with each other for solving complex graph reasoning problems.
+Extensive experiments show that GraphTeam achieves state-of-the-art results with impressive accuracy improvements.
+As the first work to address LLM-based graph reasoning problems using a multi-agent system, our workflow prioritizes simplicity, with most agents operating sequentially.
+For future work, we will consider to explore more sophisticated communications between agents, and automatically optimize the system pipeline to accommodate more tasks.
+It is also possible to fine-tune the LLMs in GraphTeam for better performance.
+Appendix A Method Details
+A.1 Problem-Solving Experience Collection
+[…]
+Algorithm 2 Problem-Solving Experience Collection
+0:  Training dataset $\mathcal{D}_{\text{train}}$, validation dataset $\mathcal{D}_{\text{valid}}$, candidate experience size $N_{exp}$ for each problem type $t\in\mathcal{T}$.
+0:  Experience knowledge base $\mathcal{K}_{exp}$.
+1:  for each problem $(q,a,t)$ from $\mathcal{D}_{\text{train}}$ do
+2:     if $|\mathcal{K}_{exp}^{t}|$ $<$ $N_{exp}$ then
+[…]
+4:        if answer equals to ground truth $a$ then
+[…]
+6:        end if
+7:     end if
+8:  end for
+9:  for each problem $(q,a,t)$ from $\mathcal{D}_{\text{valid}}$ do
+10:     for $e\in\mathcal{K}_{exp}^{t}$ do
+[…]
+12:        if answer equals to ground truth $a$ then
+13:           $\text{Utility}_{e}\leftarrow\text{Utility}_{e}+1$
+14:        end if
+15:     end for
+16:  end for
+17:  for $t\in\mathcal{T}$ do
+18:     $\mathcal{K}_{exp}^{t}\leftarrow\text{argmax}_{e\in\mathcal{K}_{exp}^{t}}\text{Utility}_{e}$
+19:  end for
+20:  $\mathcal{K}_{exp}\leftarrow\{\mathcal{K}_{exp}^{t}\}_{t\in\mathcal{T}}$
+A.2 Prompt Templates
+Below are the prompts of the five agents used in GraphTeam.
+A.2.1 Question Agent
+Here is the task: {original question}.
+As a requirement analyst, you need to extract key information from the task, specifically focusing on the following elements:
+Reformatted_Problem (str): This is the specific problem that needs to be solved.
 Graph_Type (str): This refers to the type of graph mentioned in the task, such as directed, undirected, weighted, dynamic, etc.
 If the task does not specify a graph type, it should be set to undirected.
-Input(str): This refers to the graph data or graph file required by the task.
-You must extract **all relevant information** about the graph data, including but not limited to graph representation (adjacency list, adjacency matrix, edge list, etc.), node and edge details, and any other relevant information.
-Problem(str): This is the specific problem that needs to be solved.
+Input_Data (str): This refers to the graph data or graph file required by the task.
+You must extract all relevant information about the graph data, including but not limited to graph representation (adjacency list, adjacency matrix, edge list, etc.), node and edge details, and any other relevant information.
 Output_Format (str): This is the format of the output answer mentioned in the task. If the task does not specify a format, it should be set to None.
 Note that we must ONLY discuss these key elements and do not discuss anything else! Please provide the extracted information in JSON format. Your answer must be in JSON format.
-A.2. Search Agent
-Research Assistant
-You are tasked as a assistant_role to retrieve relevant information for the task: query.
+A.2.2 Search Agent
+You are tasked as a research assistant to retrieve relevant information for the task: {reformatted_question, graph_type}.
 Use the available knowledge base and memory to find the most accurate and relevant information related to the task.
@@
 Please provide the results in a structured format.
-A.3. Coding Agent
-Graph Learning Specialist
-Based on the available search result: search_result, your goal is to write Python code that solves the specified graph learning problem using libraries like NetworkX.
+A.2.3 Coding Agent
+Based on the available search result: {konwledge of the question}, your goal is to write Python codes that solve the specified graph learning problem using libraries like NetworkX.
 If there is no search result, generate the code based on your expertise.
 ATTENTION: This is the actual problem that needs to be solved, based on the provided information. Please disregard the example above while solving this real task.
+Reformatted_Problem: reformatted_problem,
 Graph_Type: graph_type
-Input: input,
-Problem: problem,
-Output Format: output_format.
+Input_Data: input_data,
+Output_Format: output_format.
 The output code must adhere to the following format:
@@
     ‘‘‘python
-A.4. Reasoning Agent
-Graph Learning Expert
-As a assistant_role, your task is to analyze and solve a graph analysis problem.
+A.2.4 Reasoning Agent
+As a graph learning expert, your task is to analyze and solve a graph reasoning problem: {refomatted_question}.
 Despite any challenges that previous attempts may have encountered, your goal is to derive a solution leveraging your expertise in graph theory and machine learning.
-Here is the task based on the extracted information: Input: input, Problem: problem, Output Format: output_format.
+Here is the task based on the extracted information: Input_Data: input_data, Reformatted_Problem: reformatted_problem, Output_Format: output_format.
 Provide a clear, concise answer and present your solution in the following format for easy extraction:
-’###output:
-solution
-A.5. Answer Agent
-Output Format Specialist
-As a assistant_role, your role is to ensure that the generated output strictly adheres to the specified format requirements.
+A.2.5 Answer Agent
+As an assistant_role, your role is to ensure that the generated output strictly adheres to the specified format requirements.
 You will be provided with both the required output format (example format) and the actual output.
 Your task is to compare the two and determine if the actual output matches the format requirements.
-If the output does not meet the format, you must adjust the format accordingly while **preserving every part of the original content, meaning, and intent without deleting or changing any sections**.
-Important: The required format provided is an **example** that illustrates the expected structure, not the actual content or final answer.
-Do **not** alter the original meaning, sections, sentences, or intent of the actual output.
+If the output does not meet the format, you must adjust the format accordingly while preserving every part of the original content, meaning, and intent without deleting or changing any sections.
+Important: The required format provided is an example that illustrates the expected structure, not the actual content or final answer.
+Do not alter the original meaning, sections, sentences, or intent of the actual output.
 Your task is only to adjust the format if necessary.
-When adjusting the format, ensure that **no parts are removed, shortened, or reworded in a way that changes the original meaning or omits content**.
+When adjusting the format, ensure that no parts are removed, shortened, or reworded in a way that changes the original meaning or omits content.
 The adjustment should focus on structural changes such as spacing, indentation, or organization.
-Once the adjustment is made, compare the adjusted output with the original to validate that **all sections, sentences, and words are fully preserved** in the final output.
+Once the adjustment is made, compare the adjusted output with the original to validate that all sections, sentences, and words are fully preserved in the final output.
 Here are the provided details:
-- **Required format (example)**: output_format
-- **Actual output** (content to be reviewed): output
-Carefully review both and ensure that the actual output conforms to the structure of the example format, while **maintaining the original content, meaning, sections, and intent**.
-Note: Only adjust the **format** if necessary. Do **not** change the actual content, meaning, or intent of the output. If the format is already correct, no adjustment is needed.
+- Required format (example): output_format
+- Actual output (content to be reviewed): output
+Carefully review both and ensure that the actual output conforms to the structure of the example format, while maintaining the original content, meaning, sections, and intent.
+Note: Only adjust the format if necessary. Do not change the actual content, meaning, or intent of the output. If the format is already correct, no adjustment is needed.
 Your response should be in JSON format with:
-- **need_adjustment** (True/False): Indicating whether an adjustment is required.
-- **output** (str): If no adjustment is needed, this should match the original output exactly.
+- need_adjustment (True/False): Indicating whether an adjustment is required.
+- output (str): If no adjustment is needed, this should match the original output exactly.
 If an adjustment is needed, this should contain the adjusted output, with only format changes applied, without altering the content, meaning, sections, or intent.
-Appendix B GNN-AutoGL Benchmark Construction
+Appendix B Details about Datasets and Metrics
+B.1 Dataset Statistics
+We list the statistics of the six benchmarks in Table 3, including task categories, output formats, maximum graph sizes, and the number of samples in experience and test sets.
+Table 3: Statistics of six graph reasoning benchmarks.
+Task Category Output Format Graph Size Total Experiences Test Set Size
+NLGraph Macro-level, Micro-level, GNN yes/no, digits, list/set, others $\sim 10^{1}$ 80 1,000
+Talk like a Graph Basic, Macro-level yes/no, digits, list/set $\sim 10^{1}$ 80 8,000
+GraphInstruct Basic, Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 170 5,100
+GraphWiz Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{2}$ 90 3,600
+LLM4DyG Basic, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 90 2,700
+GNN-AutoGL GNN others $\sim 10^{3}$ 10 200
+B.2 GNN-AutoGL Benchmark Construction
 For the GNN-AutoGL benchmark, we use a unified template for construction.
@@
 If the code fails to run, the score is zero.
-Below is a template for constructing a problem.
-In graph theory, several fundamental tasks are crucial for various applications.
-Shortest Path algorithms, like Dijkstra’s and Bellman-Ford, help find the minimum distance between nodes in a weighted graph, playing a key role in routing and network design.
-Hamiltonian Cycle detection determines whether a cycle exists that visits each vertex exactly once, a problem common in scheduling and logistics.
-Subgraph Matching focuses on finding a smaller graph pattern within a larger graph, widely used in chemical informatics and pattern recognition.
-These tasks, along with Graph Coloring, Spanning Trees, and Maximum Flow, demonstrate the diversity and utility of analyzing graph structures.
-With the rise of Graph Neural Networks (GNNs), new ways of addressing graph-based tasks, like node classification, link prediction, and clustering, have emerged, allowing models to learn from the inherent relationships in graph-structured data.
-Task: Given the dataset name dataset, can you use Python library to train a layer number-layer GNN model name model with hidden dimension as hidden dimension, dropout as dropout and predict the accuracy ?
-A Python code example for solving GNN-AutoGL problems.
-import pkg_resources
+Below is an example of problems in this benchmark.
+Given the cora dataset, can you use autogl to train a $2$-layer $gcn$ model with hidden dimension as $[128,256]$, dropout as $0.6$ and predict the accuracy ?
+A Python code example for solving the above problem is shown below.
 import autogl
-import torch
-from autogl.module.train import Acc
-from autogl.solver.utils import get_graph_labels,
-get_graph_masks
-from autogl.datasets import build_dataset_from_name
-cora_dataset = build_dataset_from_name(’cora’)
+from autogl.datasets import
+build_dataset_from_name
+cora_dataset =
+build_dataset_from_name(‘cora’)
 encoder_hp = {
-    ’num_layers’: 2,
-    ’hidden’: [128, 256],
-    ’dropout’: 0.6,
-    ’act’: ’relu’,
-    ’eps’: ’false’
+    ‘num_layers’: 2,
+    ‘hidden’: [128, 256],
+    ‘dropout’: 0.6,
+    ‘act’: ’relu’,
+    ‘eps’: ’false’
 predicted = solver.predict_proba()
-label = get_graph_labels(cora_dataset[0])
-[get_graph_masks(cora_dataset[0], ’test’)].cpu().
-numpy()
-print(’Test accuracy: ’, Acc.evaluate(predicted,
-label))
-Appendix C Dataset Statistics
-See Table 3.
-Appendix D Experimental Results
-D.1. NLGraph
-See Table 4.
-D.2. GraphWiz
-See Table 5.
-D.3. LLM4DyG
-See Table 6.
-D.4. GraphInstruct
-See Table 7 and Table 8.
-D.5. Talk Like a Graph
-Coming soon.
+label =
+get_graph_labels(cora_dataset[0])
+[get_graph_masks(cora_dataset[0],
+‘test’)].cpu().numpy()
+print(‘Test accuracy: ’,
+Acc.evaluate(predicted, label))
+B.3 Metric Details
+Specifically, we extract the generated codes from model responses, and compare the execution results with the reference solutions to assess whether the model-generated results match the target answers.
+For the first five benchmarks Fatemi et al.
+[2024], Chen et al.
+[2024a], Zhang et al.
+[2024b], Luo et al.
+[2024], Wang et al.
+[2024a], the evaluation scores are binary based on the exact matches between execution results and target answers.
+For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0.
+Otherwise, we check whether the hyper-parameters (e.g., model name, dataset name, hidden dimension, dropout) meet the requirements.
+Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
+Appendix C Additional Experiments
+Figure 7: Large scale graph experiments on NLGraph
+Table 4: Performance comparison on hard subsets of six graph analysis benchmarks in terms of accuracy (%).
+Category Models NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
+Main GPT-4o-mini 44.9 47.9 23.2 28.9 42.6 3.3
+SOTA 45.0 93.9 11.7 70.5 33.9 -
+GraphTeam (GPT-4o-mini) 98.1 99.2 73.3 76.3 98.3 48.5
+     - question agent 69.8 70.0 73.2 39.8 53.6 14.5
+Input-Output Normalization      - answer agent 97.5 97.4 70.0 75.3 96.3 45.8
+     - documentation 97.5 98.9 69.8 75.3 94.3 46.8
+Knowledge Retrieval      - experience 42.3 98.4 64.0 62.9 70.6 4.8
+     - coding agent 19.2 64.0 38.2 38.3 51.6 -
+     - retry 40.3 95.4 71.0 61.0 85.4 44.3
+Problem Solving      - reasoning agent 97.5 98.9 69.6 66.1 95.7 46.5
+C.1 Explorations on Large-scale Graphs
+The graph sizes in existing benchmarks are relatively small.
+To validate the effectiveness of GraphTeam in solving large-scale graph reasoning problems, we use NLGraph’s problem generation tool to randomly create 20 problems for each graph scale of $10^{1}$, $10^{2}$, $10^{3}$, $10^{4}$, $10^{5}$, and $10^{6}$ nodes.
+Then GraphTeam (GPT-4o-mini) is employed to solve these generated problems.
+For graphs with at least $10^{3}$ nodes, GraphTeam is asked to load them from files via programming.
+The results, shown in Figure 7, demonstrate that our multi-agent framework can effectively handle large-scale graph reasoning tasks, maintaining 90% accuracy for graphs containing $10^{6}$ nodes.
+C.2 Ablation Study: Analysis on Hard Subsets
+To investigate GraphTeam’s performance on challenging problems, we assess its performance on the “hard” subset of each benchmark.
+For GraphWiz, we adhere to the official partition and report results on their designated hard set.
+For other benchmarks, which all have their problems categorized into different types or domains, we employ a model-based approach: we identify the three categories where previous SOTA performs poorest and select problems from these categories.
+This selection method ensures we target the most challenging aspects of each benchmark.
+Tab.
+4 summarizes the system’s performance on these high-difficulty questions.
+Notably, GraphTeam maintains high accuracy on the Talk like a Graph hard subset (99.2%), and shows significant improvements over SOTA for NLGraph (98.1% v.s.
+45.0%) and LLM4DyG (98.3% v.s.
+33.9%).
+While performance drops are observed for GraphWiz and GraphInstruct compared to the general set, GraphTeam still substantially outperforms both GPT-4o-mini and SOTA methods.
+The retry mechanism shows clear benefits for hard problems with a maximum improvement of 57.8%.
+This indicates that hard problems require more complex codes, which cannot be written correctly in a single attempt and needs multiple tries to solve effectively.
+This consistent performance across diverse hard subsets underscores GraphTeam’s robustness.
+C.3 Ablation Study: Analysis on Different Groupings
+Here we further compare with the ablated variants under different groupings.
+Figure 8: Performance with respect to different task categories.
+Group by Task Categories.
+The results are shown in Figure 8.
+For micro-level tasks, removing the reasoning agent causes a significant 21.2% drop in performance, highlighting the complexity of these problems and the limitations of the coding agent alone.
+In GNN-related tasks, while the coding agent and experience are important, the question agent plays a crucial role (38.8% impact when removed).
+This underscores the importance of effectively extracting key information from GNN problem statements.
+These findings point to areas for focused improvement, particularly in enhancing reasoning capabilities for micro-level tasks and information extraction for GNN-related problems.
+Figure 9: Performance with respect to different output formats.
+Group by Output Formats.
+As shown in Figure 9, GraphTeam demonstrates strong performance on yes/no, digits, and list/set outputs, achieving an accuracy of over 85% for each of these categories.
+However, the system’s performance drops notably for more complex output formats, with an accuracy of only 70.1%.
+This decline in accuracy for the “others” format indicates that GraphTeam, particularly its answer agent, needs enhancement in handling advanced data structures and more complex output formats.
+Improving the system’s capability to generate and structure complex outputs could significantly boost its overall effectiveness in addressing a wider range of graph analysis problems.
+C.4 Detailed Results for Each Benchmark
+In this section, we present comprehensive results for the subclasses of five benchmarks: NLGraph, Talk like a graph, GraphWiz, LLM4DyG, and GraphInstruct.
+We employ GPT-4o-mini with the original state-of-the-art (SOTA) results for these benchmarks.
+C.4.1 NLGraph
+Table 5: Performance comparison on NLGraph benchmark in terms of accuracy (%).
+Connect Cycle Topo. Sort Shortest Path Max. Flow Bipartite Graph Hamilton Path GNN Overall Average
+GPT-4o-mini 72.0 52.9 20.0 35.9 98.3 29.8 22.4 0 51.4
+Previous SOTA 99.5 96.9 63.7 60.9 17.2 57.1 39.7 94.9 79.7
+GraphTeam (GPT-4o-mini) 97.0 100.0 94.8 98.4 100.0 100.0 100.0 97.4 97.8
+As shown in Table 5, we present detailed results across subclasses of NLGraph.
+GraphTeam (GPT-4o-mini) achieves perfect scores in multiple categories (including Cycle, Maximum Flow, Bipartite Graph, and Hamilton Path), with its lowest performance being 94.8 in Topology Sort.
+These results substantially exceed the previous SOTA Ouyang et al.
+[2024] and GPT-4o-mini baseline.
+C.4.2 Talk Like a Graph
+Table 6: Performance comparison on Talk like a graph benchmark in terms of accuracy (%).
+Node Count Edge Count Edge Existence Node Degree Connected Nodes Cycle Check Shortest Path Triangle Counting Overall Average
+GPT-4o-mini 100.0 45.6 87.0 54.6 50.2 92.6 45.6 14.4 61.3
+Previous SOTA 100.0 100.0 96.1 91.7 98.0 98.0 97.2 40.5 90.2
+GraphTeam (GPT-4o-mini) 100.0 99.2 61.0 98.6 99.2 99.4 99.1 99.2 94.5
+As shown in Table 6, GraphTeam (GPT-4o-mini) achieves over 90.0% accuracy across the vast majority of subclasses.
+For the edge existence task, we encounter performance limitations due to ambiguity in problem descriptions.
+Despite relatively weak performance on this task, GraphTeam (GPT-4o-mini) maintains an average performance above 90.0%, demonstrating strong generalization capability.
+C.4.3 GraphInstruct
+Table 7: Performance comparison on GraphInstruct benchmark in terms of accuracy (%).
+Neighbor Bipatite Edge Connectivity Degree DFS Predecessor Topological Sort Connected Component
+GPT-4o-mini 98.0 89.7 90.7 92.2 98.0 89.3 47.4 59.0 39.3
+Previous SOTA 99.0 85.0 84.0 83.0 81.0 46.0 41.0 33.0 83.0
+GraphTeam (GPT-4o-mini) 100.0 99.0 99.3 100.0 100.0 95.7 98.0 98.3 94.7
+Common Neighbor Hamiltonian Path Jaccard Shortest Path Diameter PageRank Maximum Flow MST Overall Average
+GPT-4o-mini 49.0 33.7 17.0 57.3 23.0 15.3 15.0 1.0 53.7
+Previous SOTA 23.0 18.0 14.0 14.0 13.0 13.0 6.0 4.0 43.5
+GraphTeam (GPT-4o-mini) 100.0 97.7 100.0 100.0 100.0 16.3 99.3 0 88.1
+As shown in Table 7, results show that GraphTeam can solve the vast majority of GraphInstruct tasks with accuracy exceeding 90.0%, except for MST (minimum spanning tree) tasks.
+The poor performance on MST tasks is attributed to a fundamental limitation where current mainstream LLMs lack the required knowledge, resulting in poor performance across all baselines as well as our GraphTeam.
+In the future, we plan to enhance GraphTeam to address such challenges.
+Nevertheless, GraphTeam significantly outperforms previous SOTA on all other subclasses of GraphInstruct.
+C.4.4 GraphWiz
+Table 8: Performance comparison on GraphWiz benchmark in terms of accuracy (%).
+Cycle Connect Bipartite Topology Shortest Triangle Flow Hamilton Subgraph Overall Average
+GPT-4o-mini 82.0 64.8 67.3 100.0 8.5 10.3 19.3 36.3 40.3 47.6
+Previous SOTA 89.0 82.5 84.8 46.8 24.0 52.8 43.5 81.5 77.3 65.0
+GraphTeam (GPT-4o-mini) 100.0 86.5 96.3 100.0 95.0 94.0 94.0 32.5 99.3 88.6
+The results are shown in Table 8.
+Although both GPT-4o-mini and GraphTeam (GPT-4o-mini) can solve the Hamilton problem conceptually, their performance is substantially impacted by the inability to match the complex output format required by the evaluation script of GraphWiz.
+The previous SOTA model Chen et al.
+[2024a] is fine-tuned on extensive datasets, and thus able to meet the required output format.
+Nevertheless, GraphTeam achieves an overall accuracy of 88.1% on GraphWiz, demonstrating strong graph reasoning capabilities.
+C.4.5 LLM4DyG
+Table 9: Performance comparison on LLM4DyG benchmark in terms of accuracy (%).
+When Link When Connect When Tclosure Neighbor at Time Neighbor in Periods Check Tclosure Check Tpath Find Tpath Sort Edge Overall Average
+GPT-4o-mini 58.0 67.3 87.7 33.3 44.7 77.3 57.7 46.7 49.7 58.0
+Previous SOTA 47.3 64.1 63.0 45.9 20.0 66.7 59.8 80.8 35.8 53.7
+GraphTeam (GPT-4o-mini) 96.3 94.3 100.0 95.3 100.0 99.7 87.0 81.0 99.7 94.8
+As shown in Table 9, GraphTeam (GPT-4o-mini) demonstrates exceptional performance, achieving over 90.0% accuracy on multiple tasks (including 100.0% accuracy on When T-closure and Neighbor in Periods).
+Even its lowest performance, on Find Tpath, reaches 81.0% accuracy, significantly outperforming previous SOTA.